### v2

# Witnessing mass-energy equivalence with trapped atom interferometers

Abstract
We propose an experimental setup to probe the interplay between the quantum superposition principle and the gravitational time dilation arising from the mass-energy equivalence. It capitalizes on state-of-the-art atom interferometers that can keep atoms trapped in a superposition of heights in Earth’s gravitational field for exceedingly long times reaching minute-scale. Our proposal consists of adding two additional laser pulses to the existing setups that would set up a clock trapped at a superposition of heights reading a quantum superposition of relativistic proper times. We develop a method to include relativistic corrections to Bloch oscillations which describe the trapped part of the interferometer. As a result, we derive all the trajectories arising in this setup, the phases acquired by different trajectories in the interferometer, and demonstrate that the effect of superposition of proper times would manifest itself in the interference pattern in two distinct ways: as visibility modulations, and as a shift of the resonant frequency of the atom. We argue that the latter might be observable with current technology.
—Introduction. A key prediction of general relativity is that time is not a global background parameter but flows at different rates depending on the spacetime geometry, a phenomenon known as time dilation [1]. The physical effects produced by time dilation have been tested in numerous paradigmatic experiments. Pound and Rebka were the first to observe an altitude-dependent gravitational redshift in gamma rays emitted between the top and bottom of a tower [2]; Hafele and Keating directly tested special and general relativistic time dilation by comparing actual clocks at different heights and moving at different speeds [3]; while Shapiro proposed to measure the reduction of the speed of light for electromagnetic waves traveling across regions subject to a gravitational potential [4], which was observed in [5]. The above effects are fully explainable within the paradigm of classical (non-quantum) relativistic physics.
The first experiment measuring the effect of gravity on the quantum wavefunction of a single particle was performed by Colella, Overhauser, and Werner (COW) [6]. The setup considered neutrons traveling in a superposition of heights in the Earth’s gravitational field, following the basic geometry of a Mach-Zehnder-type interferometer. The neutrons, traversing arms at differnt heights in the gravitational potential acquire a relative phase. The resulting interference pattern was measured and a phase difference consistent with the Earth’s gravitational potential was deduced. These and related effects measured with atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] are thus tests of the non-relativistic gravitational effects in quantum systems.
Hence, none of the experiments performed so far probe the interplay between quantum mechanics and general relativity — the idea of how to do it was only developed in [23, 24]. These works considered interferometric experiments with a particle having an internal degree of freedom that could be treated operationally as a “quantum clock”. The internal evolution would then depend on the proper time along the paths taken through general curved spacetime [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Specifically, the authors of [24] predicted that in the proposed experiment one would observe oscillations of the visibility of the interference pattern, defined as the contrast between the constructive and destructive interference fringes (which in the non-relativistic theory in the COW-type experiments is, in principle, always maximal). The predicted periodic modulations of the visibility would be an effect that arises unambiguously from the interplay between relativistic and quantum physics. Several proposals were made to probe this effect in interferometric experiments with electrons [35] and atoms [27, 28, 36, 37, 38]. However, given the size of the interferometer required for a reasonable visibility drop [24] it is rather unlikely that this effect will be observed anytime soon.
Derivations of the dynamics of relativistic quantum particles with quantized internal mass-energy, necessary to make predictions for the above experiments, have been done in the framework of relativistic quantum mechanics in refs [39, 40, 41, 42], and starting from quantum field theory in curved background in refs [43, 44, 28]. In the low energy limit for the center of mass of the composite particles, they coincide. These results were applied to modeling the dynamics of particles in clock interferometers [35, 27, 28, 36, 45, 37] and of trapped particles [46], in particular in the context of optical clocks to explore relativistic contributions to the clock’s frequency and precision in refs [47, 48, 49], and most recently to explore associated many-body effects [50].
In this work, we propose a concrete experimental setup for witnessing time dilation effects in atom interference. Our idea is inspired by the rapid development of trapped atom interferometers [18, 19, 20, 21, 22] utilizing Bloch oscillations to keep atoms in a coherent superposition of heights in a gravitational field for increasingly long times, reaching minute-scale [22]. We propose a modification of the existing interferometers in the form of two additional “clock pulses” at the beginning and the end of the Bloch oscillations. This procedure effectively sets up a clock at a superposition of heights in the gravitational field, which means that the clock experiences a superposition of proper times. We develop a method to account for relativistic mass-energy corrections in Bloch oscillations and use it to derive all the phases acquired by different trajectories within the interferometer and analyze the interference pattern. We recover the visibility modulations predicted in [24], where fixed classical trajectories of the clocks within the interferometer were assumed, but crucially, we predict the clock’s frequency shift, making this proposal feasible.
Indeed, the proposed setup implements Ramsey spectroscopy for spatially superposed atoms, allowing for measurements of their resonant frequencies. We show that relativistic effects manifest therein as resonant-frequency shifts. We argue that this will be easier to observe than the visibility oscillations due to the enormous precision of frequency measurements [51, 52]. In particular, for the height separations reported in ref. [20], the predicted fractional frequency shift is of the order of $10^{-22}$, one order of magnitude below the current measurement precision. Revealing effects potentially detectable with current technology, our proposal constitutes a major step toward experimental tests of the interplay between quantum mechanics and general relativity.
—Setup. Let us consider a one-dimensional atom interferometry setup, Fig. 1, closely resembling configurations implemented in [18, 19, 20, 21, 22]. It involves a highly cooled atom 111in experiments [20, 21, 22] the temperature reaches a few hundreds of nanokelvins moving along the gravitational field $g$ with velocity modified by Bragg pulses and an optical lattice. We assume a two-level internal structure of the atom with the ground and excited states separated by energy $\hbar\omega_{0}$.
[…]
Figure 1: Proposed setup. All motion is restricted to one dimension (parallel to the gravitational field). Blue wavy lines represent $\pi/2$ Bragg pulses splitting the trajectory of the atom, while red wavy lines represent the $\pi/2$ clock pulses affecting only the internal state of the atom. The optical lattice (green striped pattern) is turned on exactly when the blue trajectories reach their apex and off after time $T_{\text{B}}$. Shaded dashed lines represent trajectories lost within the interferometer.
First, atoms are prepared in the ground state and launched upwards into free fall with initial velocity $v_{0}$. Next, a pair of $\pi/2$ Bragg pulses are applied separated by a time $T$. Each of them splits the atomic trajectory into two — one with unchanged velocity, and one with velocity modified by $\Delta v$. Therefore, after this sequence, the initial trajectory of each atom is split into four paths with pairwise equal velocities — one pair with velocity $v_{0}-gT$, and the other pair with velocity greater by $\Delta v$. We focus on the latter pair and assume that the track of the remaining trajectories is lost as in experiments in refs [18, 19, 20, 21, 22].
When the relevant trajectories reach the apex, namely at time $T+T^{\prime}=(v_{0}+\Delta v)/g$, the optical lattice is switched on, and atoms performs Bloch oscillations. At switching on the lattice we introduce our modification — a clock pulse at frequency $\omega$ that puts atoms in an equal superposition of the two internal states without affecting its momentum [54, 55, 56]. The optical lattice remains on for time $T_{\text{B}}$, after which the second clock pulse is applied and the lattice is switched off. The atoms fall freely for time $T^{\prime}$, after which another pair of $\pi/2$ Bragg pulses is applied to recombine the trajectory. At this stage, another two trajectories are lost — this is different than in experiments [18, 19, 20, 21, 22] where all the output ports are measured but only two interfere (a modification that we introduce for simplicity of the subsequent analysis).
In the end, one measures the probabilities $P_{g}^{(j)}$ that the atoms are found in the ground state in the output trajectory $j\in\{0,1\}$, with 0 and 1 corresponding to the red and blue trajectories in Fig. 1, respectively). Apart from the non-relativistic phase $\Delta\phi$ between the upper and lower ground state trajectories (measured in typical interferometric experiments, e.g. [20]), there will be additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired within the optical lattice by the excited state on the lower and upper trajectory, respectively, which can be measured with the proposed setup. The probability $P_{g}^{(j)}$ is given by (see Appendix A for derivation)
$\begin{split}P_{g}^{(j)}=\frac{1}{16}&\bigg{[}1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}\times\\ &\times 2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\bigg{]}.\end{split}$ (1)
One can get rid of the $\Delta\phi$-dependence by calculating the total probability $P_{g}=P_{g}^{(0)}+P_{g}^{(1)}$ of finding the atom in the ground state
$P_{g}=\frac{1}{8}\left[1-\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\right].$ (2)
This probability does not depend on the specific opening and closure of the interferometer, but only on the part confined in the optical lattice. This part is central to our considerations because the atom will effectively behave as a clock at a superposition of heights within the optical lattice. The presence of two clock pulses is the main modification to previous experimental implementations [18, 19, 20, 21, 22]. Let us notice, that without those pulses, the probability of measuring the atoms in the ground state would be constant and equal to 1/4 (recall that we lose some of the trajectories). In our setup, they oscillate between 0 and 1/4, depending on the time $T_{\text{B}}$ of the lattice hold.
—Dynamics and phases. Let us analyze the optical lattice stage of the interferometer and the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state. At this stage, the atom will Bloch oscillate at a superposition of heights separated by $\Delta z$. We adopt a simplified model of Bloch oscillations (see [57, 19] and Fig. 2), in which the atom interacts with the optical lattice only when it reaches a downward velocity $v_{\text{B}}=\hbar k/m$ (here $k$ is the wave vector of the optical-lattice laser and $m$ is the mass of the ground-state atom). Then, it instantaneously exchanges two photons with the lattice and changes its velocity. For the appropriately chosen frequency of the lattice light, the velocity at the interaction point will simply flip its sign, and the atom will oscillate at the fixed height with period $\tau_{\text{B}}=2v_{\text{B}}/g$. We will assume that the laser frequency has been chosen exactly this way. Each bounce starts and ends at the apex of a free-fall trajectory. Within the time $T_{\text{B}}$ the atom performs $N$ oscillations with the last one possibly extended (or shortened) by a time $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$ (this is because $T_{\text{B}}$ is not necessarily an integer multiple of the Bloch period $\tau_{\text{B}}$; see Fig. 2).
$N\tau_{\text{B}}$$\tau$$T_{\text{B}}$
Figure 2: Trajectory followed by the atom (at the lower or upper trajectory) in the optical lattice. We adopt a simplified model of Bloch oscillations in which the atom follows a bouncing trajectory at a fixed height. The total time of Bloch oscillations $T_{\text{B}}$ needs not be equal to an integer multiple of Bloch periods $N\tau_{\text{B}}$. Instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\pi/2,\pi/2)$.
Crucially, due to the mass difference between the excited and the ground state of the atom, the condition for bouncing at the fixed height can be satisfied only for one of the internal states for a given lattice. For instance, if we choose the lattice laser frequency so that the condition is satisfied for the ground state, the excited state will necessarily fall beyond the starting height with each oscillation. However, it can be shown (see Appendix B) that in viable experimental scenarios, the total fall is much smaller than the thermal wavelength of the atom. Therefore, employing the perturbative approach developed in [58] is justified, which here means that we can approximate that both the ground and the excited state follow the same trajectory. We further verify that this approximation holds in Appendix D.
The phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ consist of two components: laser phase $\omega T_{\text{B}}$ acquired by the excited state due to the clock pulses, and a propagation phase correction due to different masses of the ground and the excited state. To calculate the latter we follow [58] and express the Lagrangian of the excited state $L_{e}=L_{g}+\delta L$ in terms of the Lagrangian $L_{g}$ of the ground state and a small correction $\delta L$ resulting from the rest-mass difference between the ground and the excited state. If we denote this mass difference by $\delta m=\hbar\omega_{0}/c^{2}$, the Lagrangian correction is given by
$\delta L=-\delta m\left(c^{2}+gz-v^{2}/2\right).$ (3)
Then, to calculate the propagation phase correction we integrate $\delta L$ over the trajectory followed by the atom and divide by $\hbar$. The full calculation is presented in the Appendix C (followed by a non-perturbative calculation in Appendix D). It leads to the following formulae for the phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$:
$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}},\end{split}$ (4)
where $\langle v^{2}\rangle=v_{\text{B}}^{2}/3$ is the mean square speed within a single oscillation.
In principle, the mass-energy corrections to the interferometric phases are not the only corrections of the order of $1/c^{2}$ that are present in the proposed experiment — following [59] we could also include corrections coming from the higher-order expansion of the free-fall Lagrangian of the relativistic particle, and $1/c^{2}$ corrections to the laser phases coming from elastic interactions (Bragg and Bloch pulses). These depend only on the trajectory followed by the atom and thus, in the picture in which ground and excited states follow the same trajectories, affect only the interferometric phase $\Delta\phi$ leaving $\delta_{\text{k}}$ and $\delta_{\text{g}}$ unmodified. However, the $1/c^{2}$ corrections to $\Delta\phi$ are not the main focus of our considerations and we will not deal with them.
On the contrary, any corrections to the clock-pulse laser phases would affect $\delta_{\text{k}}$ and $\delta_{\text{g}}$ if they were present. Let us notice, however, that the corrections to the laser phases calculated in [59] resulted from the assumption that the atom-light interactions are always resonant and the atom interacts with photons with laboratory-frame frequencies depending on its position and velocity. For our present purposes, we assume instead that the laser frequency of the clock pulses is fixed and the non-resonant processes are possible. Then, there are no corrections to the clock-pulse laser phases because the atom interacts with photons with the same laboratory-frame frequency regardless of its height and velocity.
—Results and discussion. Let us denote $\varepsilon_{\text{k}}=\langle v^{2}\rangle/2c^{2}$ for the kinetic correction in Eq. (4) and $\varepsilon_{\text{g}}=g\Delta z/c^{2}$ for the gravitational correction, and denote by $\langle\omega_{0}\rangle=\omega_{0}\left(1+\varepsilon_{\text{k}}+\varepsilon_{\text{g}}/2\right)$ the (quantum mechanical) mean of the atomic internal frequency. Then, the probability of detecting the atom in the ground state (compare with Eq. (2)) can be written as
$P_{g}=\frac{1}{8}\left[1-\mathcal{V}(T_{\text{B}})\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)\right],$ (5)
where $\mathcal{V}(T_{\text{B}}):=\cos\left(\varepsilon_{\text{g}}\omega_{0}T_{\text{B}}/2\right)$ is the interferometric visibility. In a generic situation (away from the resonance), the visibility varies very slowly with time $T_{\text{B}}$ compared to $\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)$. Therefore, Eq. (5) describes an interference pattern with a slowly varying envelope. Apart from the visibility modulations, Eq. (5) looks like a result of a Ramsey interferometry experiment for an atom with internal transition frequency $\langle\omega_{0}\rangle$. For fixed time $T_{\text{B}}$ one can scan through different clock pulse frequencies $\omega$ to find this resonant frequency. Thus, the relativistic effects coming from mass-energy equivalence can be measured in two ways: by observing the visibility oscillations, and by measuring the fractional frequency shift
$\left\langle\frac{\delta\omega_{0}}{\omega_{0}}\right\rangle\equiv\frac{\langle\omega_{0}\rangle-\omega_{0}}{\omega_{0}}=\varepsilon_{\text{k}}+\frac{\varepsilon_{\text{g}}}{2}.$ (6)
Note that while the visibility oscillations are sensitive only to the gravitational correction $\varepsilon_{\text{g}}$, the frequency shift depends both on $\varepsilon_{\text{g}}$ and $\varepsilon_{\text{k}}$. To distinguish between the gravitational and the kinetic effect in fractional frequency shift measurements, one can modulate the separation $\Delta z$ between the upper and the lower trajectory — this will affect the gravitational correction, but not the kinetic one. The probability $P_{g}$ as a function of time is visualized in Fig. 3 for exemplary values of $\omega$, $\omega_{0}$, $\varepsilon_{\text{g}}$, and $\varepsilon_{\text{k}}$.
[…]
Figure 3: Oscillations of the (normalized) probability $P_{g}$ for exemplary values $\omega=40\pi$, $\omega_{0}=1.1\omega$, and $\varepsilon_{\text{k}}=0.01$, and two different values of the gravitational correction: $\varepsilon_{\text{g}}=0$ (in gray) and $\varepsilon_{\text{g}}=\varepsilon_{\text{k}}$ (in orange). Note that, apart from visibility modulations, the orange pattern has its frequency shifted against the gray one.
[…]
Note that the observation of the frequency shift (6) itself would not be evidence of the quantum clock reading the superposition of proper times since the frequency shift is the same as for the clock being in a statistical mixture of two heights. To reject the latter possibility one can, apart from the total probability $P_{g}$ of measuring the atom in the ground state at the output, extract from the measured data also the difference $D_{g}\equiv P_{g}^{(0)}-P_{g}^{(1)}$ given by (compare with Eq. (1))
$D_{g}=\frac{1}{8}\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right),$ (7)
where $\delta_{\text{d}}$ and $\delta_{\text{u}}$ are given in Eq. (4), and $\Delta\phi$ is given by (see Appendix C for derivation)
$\Delta\phi=-\frac{mg\Delta z}{\hbar}\left(T+2T^{\prime}+\tau\right).$ (8)
If the atom is in a superposition of heights, $D_{g}$ will oscillate with time $T_{\text{B}}$. On the other hand, for the atom following a statistical mixture of two heights, the probability difference would be equal to zero. Therefore, extracting from the interferometric output the fractional frequency shift and the oscillations of $D_{g}$ would provide evidence that the atom was reading a superposition of proper times.
It is important to emphasize that the method proposed in this work does not rely on the assumption that ground and excited states follow the same trajectories within the optical lattice. While we showed that within the required order $(1/c^{2})$ of calculating the phases acquired in the interferometer the trajectories associated with different mass-energy states can be approximated as equal, the method allows accounting for mass-energy corrections and resulting phases beyond this approximation.
Finally, let us notice that the interpretation of the experiment as a spectroscopy of a clock at the superposition of heights is possible thanks to including the laser phases and the concrete mechanism that sets up the quantum clock. These were missing in the idealized scenarios in [23, 24] and in the recent work [38] proposing a similar setup using optical tweezers. Therefore, the possibility of measuring the relativistic frequency shift in such interferometric scenarios thus far went unnoticed.
Acknowledgements.
We thank Timothy Le Pers, Navdeep Arya, Germain Tobar, and Jun Ye for helpful discussions and useful comments. The authors acknowledge Knut and Alice Wallenberg foundation through a Wallenberg Academy Fellowship No. 2021.0119. J.F. acknowledges funding from the U.S. Department of Energy, Office of Science, ASCR under Award Number DE-SC0023291 and Australian Research Council Centre of Excellence for Quantum Computation and Communication Technology (No. CE1701000012).
Appendix A Probability calculation
Here we derive the formula (1) for the probability $P_{g}^{(j)}$ of measuring the atom in the ground state in output $j\in\{0,1\}$. We denote by $\Delta\phi$ the phase difference between upper and lower ground state trajectories, and by $\delta_{\text{d}}$ and $\delta_{\text{u}}$ the additional phase acquired by the excited state on the lower and upper trajectory, respectively. The final state of the atom |$\ket{\psi_{\text{final}}}$ can be written in a generic way
|$\ket{\psi_{\text{final}}}=\frac{1}{8}\ket{g}\left(\ket{0}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]-\mathrm{i}\ket{1}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}-\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]\right)+\ldots,$ (9)
where |$\ket{0}$ and |$\ket{1}$ are the states corresponding to two relevant output trajectories, and dots at the end stand for the parts of the state irrelevant to our considerations (lost trajectories and excited state). The factor $1/8=(1/\sqrt{2})^{6}$ in front comes from six laser pulses applied to the atom (four Bragg pulses and two clock pulses). The minus signs in the bracket multiplying |$\ket{0}$ result from the fact that excited state trajectories interacted with a laser two more times than the ground state trajectories (two clock pulses) — each interaction gives rise to the factor $\mathrm{i}\mathrm{e}^{\mathrm{i}\phi_{l}}$ multiplying the corresponding trajectory (here $\phi_{l}$ is the so-called laser phase and we absorb it into phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$). The factor $-\mathrm{i}$ multiplying |$\ket{1}$ reflects the fact that trajectories leaving the interferometer in output 1 interacted with the odd number of Bragg pulses (the lower trajectory interacted once, the upper one three times), while the ones leaving in output 0 interacted with the even number (two) of Bragg pulses. Finally, the additional minus sign in the bracket multiplying |$\ket{1}$ comes from the fact that, in the case of trajectories leaving in output 1, the upper trajectory interacted with two more Bragg pulses than the lower one.
The amplitude of finding the atom in the ground state in the output $j$ is given by
$\bra{g,j}\ket{\psi_{\text{final}}}\propto\frac{1}{8}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+(-1)^{j}\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right],$ (10)
where we omitted the common phase factor $-\mathrm{i}$ present in the case of $j=1$. To calculate the probability $P_{g}^{(j)}$ we take the absolute value squared of the above amplitude and get
$P_{g}^{(j)}=\left|\bra{g,j}\ket{\psi_{\text{final}}}\right|^{2}=\frac{1}{16}\left[1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\right],$ (11)
which is exactly Eq. (1).
Appendix B Fall of the excited-state trajectory
Let us motivate the use of the perturbative method [58] by analyzing the fall of the excited-state trajectory during Bloch oscillations. We assume that the frequencies of the optical lattice lasers are chosen such that the ground state follows a bouncing trajectory at a fixed height (i.e., each interaction with the laser occurs at the same height $z_{\text{B}}$). Then, the excited-state trajectory necessarily falls with each bounce, due to its greater mass. Indeed, since we assume that the interaction with the laser occurs only when the atom reaches downward velocity $v_{\text{B}}=\hbar k/m$, and the momentum transferred to the atom in each interaction is the same for both the ground and the excited state, the excited state will move after each interaction with the upward velocity $v_{\text{B}}-\delta v_{\text{B}}$, where the velocity difference $\delta v_{\text{B}}$ is given by
$\delta v_{\text{B}}=2v_{\text{B}}\frac{\delta m}{m}.$ (12)
Since the velocity of the excited state right before each interaction is larger (in terms of the absolute value) than right after it, each subsequent interaction will occur lower by
$\delta z_{\text{B}}=\frac{v_{\text{B}}\delta v_{\text{B}}}{g}$ (13)
than the previous one meaning that the excited-state trajectory will progressively fall.
Also, the excited trajectory’s smaller starting velocity means it will reach the downward velocity $v_{\text{B}}$ faster than the ground state. Two subsequent interactions on the excited-state trajectory are separated by a time $\tau_{\text{B}}-\delta\tau_{\text{B}}$, where
$\delta\tau_{\text{B}}=\frac{\delta v_{\text{B}}}{g}.$ (14)
The spatial separation between the ground and the excited state trajectories is maximal when the excited state interacts with the laser. Then, it changes the direction of motion and meets the ground state trajectory at its interaction height $z_{\text{B}}$. Therefore, the trajectories meet repeatedly (at the interaction points of the ground state trajectory), but at each subsequent meeting point the excited-state trajectory has velocity lower by $\delta v_{\text{B}}$ than at the previous meeting point (while the ground state moves there at velocity $v_{\text{B}}$).
[…]
$\lambda_{\text{th}}=\sqrt{\frac{2\pi\hbar^{2}}{mk_{\text{\tiny{B}}}T}},$ (15)
where $T$ is the temperature. In the experiment [20] the atoms were cooled to $\sim 400\;\text{nK}$, corresponding to $\lambda_{\text{th}}\sim 10^{-7}\;\text{m}$. Therefore, the thermal wavelength is much larger than the separation between the trajectories, and it is justified to use the perturbative approach in phase calculations.
Appendix C Phases
Let us calculate the phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$ using the perturbative method developed in [58]. Each consists of two contributions: the propagation phase calculated by integrating the free-fall Lagrangian over the trajectory followed by the atom, and the laser phase resulting from atom-light interactions.
The free-fall Lagrangian of a relativistic particle with the internal two-level structure incorporating the mass-energy equivalence can be expanded to the order $1/c^{2}$ as follows:
$\begin{split}L=-\hat{m}c^{2}-\hat{m}\left(gz-\frac{1}{2}v^{2}\right)-\frac{\hat{m}}{c^{2}}\left(\frac{1}{2}g^{2}z^{2}-\frac{1}{8}v^{4}+\frac{3}{2}gzv^{2}\right)+\mathcal{O}\left(\frac{1}{c^{4}}\right).\end{split}$ (16)
Here $\hat{m}$ is the mass operator returning $m$ for the ground state, and $m+\delta m$ (where $\delta m=\hbar\omega_{0}/c^{2}$) for the excited state of the atom. At this point, we distinguish the $1/c^{2}$ correction to the excited-state Lagrangian coming from the mass-energy equivalence
$\delta L^{\text{(1)}}=-\delta m\left(gz-v^{2}/2\right),$ (17)
and the correction to the (ground- and excited-state) Lagrangian coming from the expansion of the relativistic-particle Lagrangian up to $1/c^{2}$ order
$\delta L^{\text{(2)}}=-\frac{m}{c^{2}}\left(\frac{1}{2}g^{2}z^{2}-\frac{1}{8}v^{4}+\frac{3}{2}gzv^{2}\right).$ (18)
Here we replaced $\hat{m}$ by $m$ because $\delta m$ is of the order $1/c^{2}$ and would contribute in higher order to the above expression. Note that for the scenario considered in this work, based on numbers from [20], $\delta L^{(2)}\ll\delta L^{(1)}$. This is because the internal energy of the atomic transition ($\sim 1\;\text{eV}$) is much larger than the kinetic energy of the atom or the gravitational energy difference between the trajectories (both $\sim 10^{-10}\;\text{eV}$).
Note also that the correction $\delta L^{(2)}$ and the corrections to the laser phases derived in [59] depend only on the trajectory, not on the internal state. Therefore, in the picture in which both the ground and the excited state follow the same trajectory, they will contribute only to the phase $\Delta\phi$, not to $\delta_{\text{d}}$ and $\delta_{\text{u}}$. We are not interested in $1/c^{2}$ corrections to $\Delta\phi$, hence we will omit them in the subsequent analysis and focus only on the corrections coming from $\delta L^{(1)}$.
We begin with the calculation of $\Delta\phi$. To calculate the propagation phase contribution we divide the trajectory into freely falling pieces between points of interaction with the lasers and calculate the propagation phase by integrating the freely falling Lagrangian
$L_{g}=-m\left(c^{2}+gz-v^{2}/2\right)$ (19)
along the trajectory. More precisely, for the free fall starting at time $t_{\text{A}}$ and ending at $t_{\text{B}}$, the propagation phase is given by
$\begin{split}\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=&\frac{1}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}tL_{g}=-\frac{m}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}t\left(c^{2}+gz-v^{2}/2\right)\\ =&-\frac{m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right],\end{split}$ (20)
where $z_{\text{A}}$ and $v_{\text{A}}$ are the atom’s position and velocity, respectively, at time $t_{\text{A}}$. Denote by $t_{j}$ with $j\in\{1,2,3,4\}$ the times at which consecutive Bragg pulses are applied, and by $t_{\text{i}}$ and $t_{\text{f}}$ the times of the initial and final clock pulses, respectively. Notice that to match the notation from the main text, the times of application of particular pulses must satisfy
$t_{2}-t_{1}=t_{4}-t_{3}=T,\qquad t_{\text{i}}-t_{2}=t_{3}-t_{\text{f}}=T^{\prime},\qquad t_{\text{f}}-t_{\text{i}}=T_{\text{B}}.$ (21)
Let us further denote the quantities corresponding to the lower and upper trajectories by superscripts (d) and (u), respectively, and by $\Delta\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=\phi_{\text{p}}^{\text{(u)}}(t_{\text{A}}\to t_{\text{B}})-\phi_{\text{p}}^{\text{(d)}}(t_{\text{A}}\to t_{\text{B}})$ the propagation phase difference between those two trajectories acquired between $t_{\text{A}}$ and $t_{\text{B}}$. Let us consider one by one the consecutive stages of the interferometer and calculate the corresponding phase difference using (20).
  • $t_{1}\to t_{2}$:
    At time $t_{1}$ both trajectories are at the same height, but the upper one starts with a velocity greater by $\Delta v$ than the lower one. Denote the initial velocity of the lower trajectory by $v_{1}$. Then, the relative propagation phase acquired between $t_{1}$ and $t_{2}$ reads
    $\Delta\phi_{\text{p}}(t_{1}\to t_{2})=\frac{m}{\hbar}\Delta vT\left(v_{1}+\frac{1}{2}\Delta v-gT\right).$ (22)
  • $t_{2}\to t_{\text{i}}$:
    Both trajectories start with the same velocity, but the upper one is higher by $\Delta vT$. The relative propagation phase is given by
    $\Delta\phi_{\text{p}}(t_{2}\to t_{\text{i}})=-\frac{m}{\hbar}g\Delta vTT^{\prime}.$ (23)
  • $t_{\text{i}}\to t_{\text{f}}$:
    We assume that both trajectories oscillate with the same frequency, and start simultaneously with zero initial velocity and maximal height. Therefore, the trajectories are constantly separated by $\Delta vT$ and have the same velocities all the time. The relative propagation phase equals
    $\Delta\phi_{\text{p}}(t_{\text{i}}\to t_{\text{f}})=-\frac{m}{\hbar}g\Delta vTT_{\text{B}}.$ (24)
  • $t_{\text{f}}\to t_{3}$:
    Similar to the previous two points, the velocity on both trajectories is the same, and they are separated by $\Delta vT$. The propagation phase difference reads
    $\Delta\phi_{\text{p}}(t_{\text{f}}\to t_{\text{3}})=-\frac{m}{\hbar}g\Delta vTT^{\prime}.$ (25)
  • $t_{3}\to t_{4}$:
    The trajectories start separated by $\Delta vT$, and the upper one has a velocity smaller by $\Delta v$. Denote the initial velocity of the lower trajectory by $v_{3}$. The relative propagation phase is then given by
    $\Delta\phi_{\text{p}}(t_{3}\to t_{4})=\frac{m}{\hbar}\Delta vT\left(-v_{3}+\frac{1}{2}\Delta v\right).$ (26)
To calculate the total propagation phase difference $\Delta\phi_{\text{p}}$ we sum up all the contributions, which gives
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta vT\left[v_{1}+\Delta v-v_{3}-g\left(T+2T^{\prime}+T_{\text{B}}\right)\right].$ (27)
Finally, let us note that, because of the requirement that at $t_{\text{i}}$ the velocity on both trajectories is equal to zero, the following relation holds
$v_{1}+\Delta v=g(T+T^{\prime}).$ (28)
On the other hand, $v_{3}$ is the velocity that the atom reaches after time $T^{\prime}$ of a free fall after leaving the optical lattice. Let us notice that the atom does not leave the optical lattice with zero velocity (see Fig. 2), but rather with velocity $-g\tau$. Here $\tau\in\{-\tau_{\text{B}}/2,\tau_{\text{B}}/2\}$ is given by
$T_{\text{B}}=N\tau_{\text{B}}+\tau,$ (29)
where $N=\lfloor T_{\text{B}}/\tau_{\text{B}}+1/2\rfloor$ is the number of interactions between the optical lattice and the atom. Therefore, we have
$v_{3}=-g(\tau+T^{\prime})$ (30)
and we can rewrite Eq. (27) (introducing $\Delta z=\Delta vT$) as follows:
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta zg(\tau-T_{\text{B}})=-\frac{m}{\hbar}g\Delta zN\tau_{\text{B}}.$ (31)
Now, let us focus on the laser phase contribution to the phase $\Delta\phi$. As described in [58], when the atom absorbs (emits) a photon with frequency $\omega$ and wave vector $k$, it acquires a laser phase
$\phi_{\text{l}}(t,z)=\mp\mathrm{i}\omega t\pm kz,$ (32)
where the upper (lower) sign corresponds to photon absorption (emission), and $t$ and $z$ are the time and position of the interaction, respectively. In the case of Bragg pulses, as well as in the optical lattice, the photon absorption is immediately followed by the emission of a photon with the same frequency and opposite wave vector. Therefore, the laser phase acquired at each such interaction point is equal to $\phi_{\text{l}}(t,z)=\phi_{\text{l}}(z)=\pm 2kz$ with the plus (minus) sign corresponding to the atom receiving upward (downward) momentum.
Denote by $z_{j}$ with $j\in\{1,2,3,4\}$ the interaction heights with four consecutive Bragg pulses ($z_{1}$ and $z_{3}$ correspond to the upper trajectory, while $z_{2}$ and $z_{4}$ to the lower one), and by $z_{\text{d}}$ and $z_{\text{u}}$ the heights of interactions of the lower and upper trajectories, respectively, within the optical lattice. Note also that the wave vector $k_{\text{Bragg}}$ of the pulses is related to the velocity change $\Delta v$ by
$2\hbar k_{\text{Bragg}}=m\Delta v,\qquad\Longrightarrow\qquad k_{\text{Bragg}}=\frac{m\Delta v}{2\hbar},$ (33)
and the wave vector $k_{\text{Bloch}}$ of the optical lattice laser is related to the Bloch period $\tau_{\text{B}}$ by
$2\hbar k_{\text{Bloch}}=mg\tau_{\text{B}},\qquad\Longrightarrow\qquad k_{\text{Bloch}}=\frac{mg\tau_{\text{B}}}{2\hbar}.$ (34)
The total laser phase difference (between the upper and lower trajectory) is given by
$\Delta\phi_{\text{l}}=2k_{\text{Bragg}}(z_{1}-z_{2}-z_{3}+z_{4})+2Nk_{\text{Bloch}}(z_{\text{u}}-z_{\text{d}})$ (35)
Let us notice the following relations between particular heights:
$z_{2}-z_{1}=v_{1}T-\frac{1}{2}gT^{2},\qquad z_{4}-z_{3}=(v_{3}-\Delta v)T-\frac{1}{2}gT^{2},\qquad z_{\text{u}}-z_{\text{d}}=\Delta z,$ (36)
where $v_{1}$ and $v_{3}$ are given by (28) and (30). Hence, we can write the total laser phase difference as
$\Delta\phi_{\text{l}}=\frac{m}{\hbar}\left[-g\Delta z(T+2T^{\prime}+\tau)+g\Delta zN\tau_{\text{B}}\right].$ (37)
The total phase difference $\Delta\phi$ between the upper and the lower (ground state) trajectories is calculated by summing up the total propagation phase difference and the total laser phase difference. This gives
$\Delta\phi=\Delta\phi_{\text{p}}+\Delta\phi_{\text{l}}=-\frac{mg\Delta z}{\hbar}(T+2T^{\prime}+\tau).$ (38)
Now, let us calculate the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state on the lower and upper trajectory, respectively. We need to calculate the additional laser phase coming from the clock pulses and the correction to the propagation phase due to the mass-energy equivalence. Starting with the laser phase, we assume that the clock pulses change only the atom’s internal state, but do not affect its motion. This can be achieved by driving two-photon transitions with both photons having the same frequency $\omega/2$ and opposite momenta. The corresponding laser phase is equal to $\mp\omega t$ with the minus (plus) sign corresponding to the two-photon absorption (emission). Note that this phase is independent of the position $z$ of the interaction (thus, it is the same for the lower and upper trajectory). Since the excited state trajectories absorb the photons at $t_{\text{i}}$ and emit at $t_{\text{f}}$, they acquire the laser phase $\omega(t_{\text{f}}-t_{\text{i}})=\omega T_{\text{B}}$.
To calculate the propagation phase correction, we employ the perturbative method developed in [58] and integrate the mass-energy correction to the free-fall Lagrangian
$\delta L=-\delta m\left(c^{2}+gz-v^{2}/2\right)$ (39)
over the ground state trajectory on the corresponding height. For the free-fall trajectory starting at $t_{\text{A}}$ and ending at $t_{\text{B}}$ the correction $\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})$ to the propagation phase reads (compare with (20))
$\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=-\frac{\delta m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right].$ (40)
For a full single oscillation in the optical lattice, the propagation phase correction $\delta_{\text{p,1}}$ is given by
$\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (41)
where $z_{\text{B}}$ and $v_{\text{B}}$ are the atom’s position and velocity at the interaction point, and $\tau_{\text{B}}$ is the period of the oscillation. We assume that the total number $N$ of Bloch oscillations is large, and neglect the correction coming from the fact that the total time $T_{\text{B}}$ of Bloch oscillations is not necessarily equal to the integer multiple of Bloch periods (instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$; however, since $\tau\ll N\tau_{\text{B}}$, we will assume $\tau\approx 0$). Then, the total propagation phase correction is given by
$\delta_{\text{p,tot}}=N\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)N\tau_{\text{B}}\approx\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)T_{\text{B}}.$ (42)
Without loss of generality, we can assume that the interaction height corresponding to the lower trajectory $z_{\text{B}}^{\text{(d)}}=0$, while the interaction height of the upper one equals $z_{\text{B}}^{\text{(u)}}=\Delta z$. We introduce the mean velocity (within a single oscillation) $\langle v^{2}\rangle\equiv v_{\text{B}}^{2}/3$, and note that $\delta m=\omega_{0}/c^{2}$. Combining the propagation phase with the laser phase corrections, we get
$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}}.\end{split}$ (43)
Appendix D Non-perturbative calculations
As a consistency check let us compute the phases accumulated during the Bloch oscillations non-perturbatively, namely for the trajectories described in Appendix B, and make sure that they agree with the perturbative calculation from Appendix C. In the following, we restrict all the calculations to the terms up to order $1/c^{2}$ neglecting the higher-order terms. This is consistent with the perturbative approach, where we considered only $1/c^{2}$ corrections.
Since the trajectories of the ground and excited states no longer coincide (the excited trajectory falls), the phase difference between them will gain contribution from three different sources: the propagation phase, the laser phase, and the separation phase. The first two are defined in the same way as in the previous (perturbative) considerations, and the third one is given by
$\delta_{\text{s}}(t)\equiv\frac{\bar{p}(t)\Delta z(t)}{\hbar},$ (44)
where $\bar{p}(t)$ is the mean momentum of two trajectories at time $t$ (the oscillation starts at $t=0$), and $\Delta z(t)$ is the separation between them (note that this has nothing to do with the separation $\Delta z$ between the upper and lower trajectory).
Since the phase acquired due to the interaction with the clock pulses is position-independent, it will not change in the non-perturbative analysis. Therefore, we consider only the phases acquired during the Bloch oscillations within the optical lattice. The trajectory followed by both trajectories has been described above. We analyze only the time between the two first interactions on the ground state trajectory (at some arbitrary height $z_{\text{B}}$) and show that the phase difference between the ground and the excited state trajectories agrees with the perturbative calculations. It is straightforward that it will agree with the perturbative calculations also at later times.
The propagation phases acquired on both trajectories in the first part of the oscillation (before the second interaction of the excited state) can be calculated using the general formula for the propagation phase in free fall:
$\phi_{\text{p}}^{\text{(g/e)}}(t)=\frac{m^{\text{(g/e)}}}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{0}^{2}-gz_{0}\right)t-v_{0}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (45)
(this is just Eq. (20) with $t_{\text{A}}\to 0$ and $t_{\text{B}}\to t$) where $m^{\text{(g)}}=m$ and $m^{\text{(e)}}=m+\delta m$ are the masses of the atom in the ground and excited state, respectively, $v_{0}$ is the initial velocity, and $z_{0}$ is the initial height. For the ground state, we replace $v_{0}$ by $v_{\text{B}}$, while for the excited state by $v_{\text{B}}-\delta v_{\text{B}}$. The initial height for both states is the same and equal to $z_{\text{B}}$.
The difference in propagation phases at time $t$ is then given by
$\delta_{\text{p}}(t)=\phi_{\text{p}}^{\text{(e)}}(t)-\phi_{\text{p}}^{\text{(g)}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right]-\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t$ (46)
(recall that we neglect the terms higher-order than $1/c^{2}$). Let us notice that, since $\delta v_{\text{B}}t$ is just the separation $\Delta z(t)$ between the trajectories, the second term in the above formula is equal in value but with the opposite sign to the separation phase
$\delta_{\text{s}}(t)=\frac{\bar{p}(t)\delta v_{\text{B}}t}{\hbar}=\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t,$ (47)
where $\bar{p}(t)=m(v_{\text{B}}-gt)+\mathcal{O}(1/c^{2})$. Therefore, the total phase difference for $t<\tau_{\text{B}}-\delta v_{\text{B}}/g$ is given by
$\delta_{\text{p}}(t)+\delta_{\text{s}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (48)
in agreement with the perturbative results (indeed, this is just the propagation phase correction calculated in the perturbative approach).
At the time $\tau_{\text{B}}-\delta v_{\text{B}}/g$ the interaction of the atom on the excited trajectory with the laser occurs, and it acquires a laser phase
$\phi_{\text{l}}^{\text{(e)}}=\frac{m}{\hbar}2v_{\text{B}}(z_{\text{B}}-\delta z_{\text{B}})=\phi_{\text{l}}^{\text{(g)}}-\frac{m}{\hbar}\delta v_{\text{B}}\tau_{\text{B}}+\mathcal{O}(\delta m^{2}),$ (49)
where $\phi_{\text{l}}^{\text{(g)}}$ is the laser phase acquired by the ground state trajectory at time $\tau_{\text{B}}$. Let us notice that right before the second interaction of the excited state the separation phase is equal to
$\delta_{\text{s}}(\tau_{\text{B}}-\Delta v/g-\varepsilon)=-\frac{m}{\hbar}v_{\text{B}}\delta v_{\text{B}}t,$ (50)
while right after the interaction it is $\sim\mathcal{O}(1/c^{4})$, because the mean momentum is then $\sim\mathcal{O}(1/c^{2})$. This situation (mean momentum $\sim\mathcal{O}(1/c^{2})$ and thus negligible value of separation phase) persists until the second interaction on the ground state trajectory, at which time both trajectories are at the same height, but the excited one has the velocity $v_{\text{B}}-2\delta v_{\text{B}}$. At this point, we can repeat the whole analysis changing everywhere $\delta v_{\text{B}}\to 2\delta v_{\text{B}}$. Since the propagation phase difference acquired between the second interaction on the excited and on the ground trajectories is also $\sim\mathcal{O}(\delta m^{2})$, the total phase difference after the first oscillation is equal to
$\delta=\delta_{\text{p}}+\delta_{\text{s}}+\delta_{\text{l}}=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)\tau_{\text{B}}-v_{\text{B}}g\tau_{\text{B}}^{2}+\frac{1}{3}g^{2}\tau_{\text{B}}^{3}\right]=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (51)
which is equal to the propagation phase correction calculated for a single oscillation in the perturbative approach (compare with Eq. (41)).

### v3

@@
 Abstract
-We propose an experimental setup to probe the interplay between the quantum superposition principle and the gravitational time dilation arising from the mass-energy equivalence.
-It capitalizes on state-of-the-art atom interferometers that can keep atoms trapped in a superposition of heights in Earth’s gravitational field for exceedingly long times reaching minute-scale.
-Our proposal consists of adding two additional laser pulses to the existing setups that would set up a clock trapped at a superposition of heights reading a quantum superposition of relativistic proper times.
-We develop a method to include relativistic corrections to Bloch oscillations which describe the trapped part of the interferometer.
-As a result, we derive all the trajectories arising in this setup, the phases acquired by different trajectories in the interferometer, and demonstrate that the effect of superposition of proper times would manifest itself in the interference pattern in two distinct ways: as visibility modulations, and as a shift of the resonant frequency of the atom.
+We propose an experimental setup to probe the interplay between the quantum superposition principle and gravitational time dilation arising from the mass-energy equivalence.
+It capitalizes on state-of-the-art atom interferometers that can keep atoms trapped in a superposition of heights in Earth’s gravitational field for exceedingly long times, reaching the minute scale.
+Our proposal consists of adding two additional laser pulses to the existing experiments that would set up a clock trapped at a superposition of heights, reading a quantum superposition of relativistic proper times.
+We develop a method to include relativistic corrections to Bloch oscillations, which describe the trapped part of the interferometer.
+We derive the trajectories and corresponding phases acquired in each arm of the interferometer.
+We then show that a superposition of proper times manifests in the interference pattern in two ways: visibility modulations and a shift of the atom’s resonant frequency.
 We argue that the latter might be observable with current technology.
@@
 A key prediction of general relativity is that time is not a global background parameter but flows at different rates depending on the spacetime geometry, a phenomenon known as time dilation [1].
-The physical effects produced by time dilation have been tested in numerous paradigmatic experiments.
-Pound and Rebka were the first to observe an altitude-dependent gravitational redshift in gamma rays emitted between the top and bottom of a tower [2]; Hafele and Keating directly tested special and general relativistic time dilation by comparing actual clocks at different heights and moving at different speeds [3]; while Shapiro proposed to measure the reduction of the speed of light for electromagnetic waves traveling across regions subject to a gravitational potential [4], which was observed in [5].
-The above effects are fully explainable within the paradigm of classical (non-quantum) relativistic physics.
-The first experiment measuring the effect of gravity on the quantum wavefunction of a single particle was performed by Colella, Overhauser, and Werner (COW) [6].
-The setup considered neutrons traveling in a superposition of heights in the Earth’s gravitational field, following the basic geometry of a Mach-Zehnder-type interferometer.
-The neutrons, traversing arms at differnt heights in the gravitational potential acquire a relative phase.
-The resulting interference pattern was measured and a phase difference consistent with the Earth’s gravitational potential was deduced.
-These and related effects measured with atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] are thus tests of the non-relativistic gravitational effects in quantum systems.
-Hence, none of the experiments performed so far probe the interplay between quantum mechanics and general relativity — the idea of how to do it was only developed in [23, 24].
-These works considered interferometric experiments with a particle having an internal degree of freedom that could be treated operationally as a “quantum clock”.
-The internal evolution would then depend on the proper time along the paths taken through general curved spacetime [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
-Specifically, the authors of [24] predicted that in the proposed experiment one would observe oscillations of the visibility of the interference pattern, defined as the contrast between the constructive and destructive interference fringes (which in the non-relativistic theory in the COW-type experiments is, in principle, always maximal).
+The physical effects produced by time dilation have been tested in numerous paradigmatic experiments: Pound and Rebka were the first to observe an altitude-dependent gravitational redshift in gamma rays emitted between the top and bottom of a tower [2]; Hafele and Keating directly tested special and general relativistic time dilation by comparing atomic clocks at different heights and moving at different speeds [3]; and Shapiro proposed to measure the reduction in the speed of light for electromagnetic waves traveling across regions subject to a gravitational potential [4], an effect later observed in [5].
+These effects are fully explainable in classical relativistic physics.
+The first experiment to measure the effect of gravity on the quantum wavefunction of a single particle was performed by Colella, Overhauser, and Werner (COW) [6].
+In this setup, neutrons travel in a superposition of heights in Earth’s gravitational field, following the geometry of a Mach-Zehnder-type interferometer.
+The neutrons, traversing arms at different heights in the gravitational potential, acquire a relative phase.
+The resulting interference pattern was measured, yielding a phase shift consistent with Earth’s gravitational potential.
+These experiments, and related measurements with atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], thus test non-relativistic gravitational effects in quantum systems.
+The proposal to unambiguously probe the interplay between quantum mechanics and general relativity was developed in [23, 24].
+These works considered interferometric experiments with particles having an internal degree of freedom that can be treated operationally as a “quantum clock”.
+The internal evolution would then depend on the proper time along the paths taken through curved spacetime [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
+Specifically, the authors of [24] predicted that the experiment would exhibit oscillations in the interference visibility — defined as the contrast between constructive and destructive fringes — which, in non-relativistic COW-type treatments, is always maximal in principle.
 The predicted periodic modulations of the visibility would be an effect that arises unambiguously from the interplay between relativistic and quantum physics.
-Several proposals were made to probe this effect in interferometric experiments with electrons [35] and atoms [27, 28, 36, 37, 38].
-However, given the size of the interferometer required for a reasonable visibility drop [24] it is rather unlikely that this effect will be observed anytime soon.
-Derivations of the dynamics of relativistic quantum particles with quantized internal mass-energy, necessary to make predictions for the above experiments, have been done in the framework of relativistic quantum mechanics in refs [39, 40, 41, 42], and starting from quantum field theory in curved background in refs [43, 44, 28].
-In the low energy limit for the center of mass of the composite particles, they coincide.
-These results were applied to modeling the dynamics of particles in clock interferometers [35, 27, 28, 36, 45, 37] and of trapped particles [46], in particular in the context of optical clocks to explore relativistic contributions to the clock’s frequency and precision in refs [47, 48, 49], and most recently to explore associated many-body effects [50].
-In this work, we propose a concrete experimental setup for witnessing time dilation effects in atom interference.
-Our idea is inspired by the rapid development of trapped atom interferometers [18, 19, 20, 21, 22] utilizing Bloch oscillations to keep atoms in a coherent superposition of heights in a gravitational field for increasingly long times, reaching minute-scale [22].
-We propose a modification of the existing interferometers in the form of two additional “clock pulses” at the beginning and the end of the Bloch oscillations.
-This procedure effectively sets up a clock at a superposition of heights in the gravitational field, which means that the clock experiences a superposition of proper times.
-We develop a method to account for relativistic mass-energy corrections in Bloch oscillations and use it to derive all the phases acquired by different trajectories within the interferometer and analyze the interference pattern.
-We recover the visibility modulations predicted in [24], where fixed classical trajectories of the clocks within the interferometer were assumed, but crucially, we predict the clock’s frequency shift, making this proposal feasible.
+Several proposals have been made to investigate this effect in interferometric experiments with electrons [35], and with atoms [27, 28, 36, 37, 38].
+However, given the size of the interferometer required for a reasonable visibility drop [24], it is rather unlikely that this effect will be observed in the near future.
+Derivations of the dynamics of relativistic quantum particles with quantized internal mass-energy — necessary for predicting the above experiments — have been performed both within relativistic quantum mechanics [39, 40, 41, 42] and via quantum field theory on curved backgrounds [43, 44, 28].
+In the low-energy limit for the center-of-mass motion of the composite particles, these two approaches coincide.
+These results have been applied to model particle dynamics in clock interferometers [35, 27, 28, 36, 45, 37] and in trapped-particle setups [46, 47].
+In particular, they have been used in optical-clock studies to explore relativistic contributions to the clock’s frequency and precision in  [48, 49, 50], and, most recently, to explore associated many-body effects [51].
+In this work, we propose a concrete experimental setup for witnessing time dilation effects in atom interferometry.
+Our idea is inspired by the rapid development of trapped atom interferometers [18, 19, 20, 21, 22] utilizing Bloch oscillations to keep atoms in a coherent superposition of heights in a gravitational field for increasingly long times, reaching the minute scale [22].
+We propose a modification of the existing interferometers by adding two “clock pulses” at the beginning and end of the Bloch oscillations, effectively placing a clock in a superposition of heights and thus of proper times.
+We develop a method to include relativistic mass–energy corrections in the Bloch oscillations and use it to derive all phases acquired along the interferometer trajectories.
+We then analyze the resulting interference pattern.
+We recover the visibility modulations predicted in [24], which assumed fixed classical trajectories of the clocks, and crucially predict a clock frequency shift that renders this proposal experimentally feasible.
 Indeed, the proposed setup implements Ramsey spectroscopy for spatially superposed atoms, allowing for measurements of their resonant frequencies.
 We show that relativistic effects manifest therein as resonant-frequency shifts.
-We argue that this will be easier to observe than the visibility oscillations due to the enormous precision of frequency measurements [51, 52].
-In particular, for the height separations reported in ref.
-[20], the predicted fractional frequency shift is of the order of $10^{-22}$, one order of magnitude below the current measurement precision.
-Revealing effects potentially detectable with current technology, our proposal constitutes a major step toward experimental tests of the interplay between quantum mechanics and general relativity.
+We argue that this will be easier to observe than the visibility oscillations thanks to the extreme precision of frequency measurements [52, 53].
+For the height separations reported in ref.
+[20], the predicted fractional frequency shift is of order $10^{-22}$ — one order of magnitude smaller than current experimental precision.
+By revealing effects potentially detectable with existing technology, our proposal represents a major step toward experimental tests of the interplay between quantum mechanics and general relativity.
 —Setup.
@@
 1, closely resembling configurations implemented in [18, 19, 20, 21, 22].
-It involves a highly cooled atom 111in experiments [20, 21, 22] the temperature reaches a few hundreds of nanokelvins moving along the gravitational field $g$ with velocity modified by Bragg pulses and an optical lattice.
-We assume a two-level internal structure of the atom with the ground and excited states separated by energy $\hbar\omega_{0}$.
-[…]
+It involves highly cooled atoms [^1] In experiments [20, 21, 22], the temperature reaches a few hundred nanokelvins.
+moving along the gravitational field $g$, with their velocity manipulated by Bragg pulses and an optical lattice.
+We assume that atoms have a two-level internal structure, with ground and excited states separated by energy $\hbar\omega_{0}$.
+$_{\rm B}TT$$^{\prime}TT$$T^{\prime}$
 Figure 1: Proposed setup.
@@
 Blue wavy lines represent $\pi/2$ Bragg pulses splitting the trajectory of the atom, while red wavy lines represent the $\pi/2$ clock pulses affecting only the internal state of the atom.
-The optical lattice (green striped pattern) is turned on exactly when the blue trajectories reach their apex and off after time $T_{\text{B}}$.
+The optical lattice (green striped pattern) is turned on exactly when the blue trajectories reach their apex and off after time $T_{\rm B}$.
 Shaded dashed lines represent trajectories lost within the interferometer.
-First, atoms are prepared in the ground state and launched upwards into free fall with initial velocity $v_{0}$.
-Next, a pair of $\pi/2$ Bragg pulses are applied separated by a time $T$.
-Each of them splits the atomic trajectory into two — one with unchanged velocity, and one with velocity modified by $\Delta v$.
-Therefore, after this sequence, the initial trajectory of each atom is split into four paths with pairwise equal velocities — one pair with velocity $v_{0}-gT$, and the other pair with velocity greater by $\Delta v$.
-We focus on the latter pair and assume that the track of the remaining trajectories is lost as in experiments in refs [18, 19, 20, 21, 22].
-When the relevant trajectories reach the apex, namely at time $T+T^{\prime}=(v_{0}+\Delta v)/g$, the optical lattice is switched on, and atoms performs Bloch oscillations.
-At switching on the lattice we introduce our modification — a clock pulse at frequency $\omega$ that puts atoms in an equal superposition of the two internal states without affecting its momentum [54, 55, 56].
-The optical lattice remains on for time $T_{\text{B}}$, after which the second clock pulse is applied and the lattice is switched off.
-The atoms fall freely for time $T^{\prime}$, after which another pair of $\pi/2$ Bragg pulses is applied to recombine the trajectory.
-At this stage, another two trajectories are lost — this is different than in experiments [18, 19, 20, 21, 22] where all the output ports are measured but only two interfere (a modification that we introduce for simplicity of the subsequent analysis).
-In the end, one measures the probabilities $P_{g}^{(j)}$ that the atoms are found in the ground state in the output trajectory $j\in\{0,1\}$, with 0 and 1 corresponding to the red and blue trajectories in Fig.
+Atoms are prepared in their ground state and launched upward with initial velocity $v_{0}$, entering free fall under gravity.
+Then, two $\pi/2$ Bragg pulses are applied, separated by an interval $T$.
+Each pulse splits the atomic trajectory into two: one retains its velocity, while the other is shifted by $\Delta v$.
+Consequently, each initial trajectory splits into four paths with two distinct velocities: two with $v_{0}-gT$ and two with $v_{0}-gT+\Delta v$.
+We focus on the latter two paths, assuming that the others are lost, as in [18, 19, 20, 21, 22].
+When the relevant trajectories reach their apex at time $T+T^{\prime}=(v_{0}+\Delta v)/g$, the optical lattice is switched on, and atoms perform Bloch oscillations.
+When the lattice is switched on, we introduce our modification—a clock pulse at frequency $\omega$ that places atoms into an equal superposition of the two internal states without affecting their momentum [54, 55, 56, 57, 58, 59, 60].
+The optical lattice remains on for a duration $T_{\mathrm{B}}$, after which the second clock pulse is applied and the lattice is turned off.
+The atoms then fall freely for time $T^{\prime}$, after which another pair of $\pi/2$ Bragg pulses (separated by an interval $T$) is applied to recombine the trajectories.
+At this stage, two additional trajectories are lost, unlike in [18, 19, 20, 21, 22], where all output ports are measured, although only two interfere (a simplification introduced here for clarity in the subsequent analysis).
+In the end, one measures the probabilities $P_{g}^{(j)}$ that the atoms are found in the ground state in output trajectory $j\in\{0,1\}$ (with 0 and 1 corresponding to the red and blue trajectories in Fig.
 1, respectively).
-Apart from the non-relativistic phase $\Delta\phi$ between the upper and lower ground state trajectories (measured in typical interferometric experiments, e.g.
-[20]), there will be additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired within the optical lattice by the excited state on the lower and upper trajectory, respectively, which can be measured with the proposed setup.
-The probability $P_{g}^{(j)}$ is given by (see Appendix A for derivation)
-$\begin{split}P_{g}^{(j)}=\frac{1}{16}&\bigg{[}1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}\times\\ &\times 2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\bigg{]}.\end{split}$ (1)
-One can get rid of the $\Delta\phi$-dependence by calculating the total probability $P_{g}=P_{g}^{(0)}+P_{g}^{(1)}$ of finding the atom in the ground state
-$P_{g}=\frac{1}{8}\left[1-\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\right].$ (2)
-This probability does not depend on the specific opening and closure of the interferometer, but only on the part confined in the optical lattice.
-This part is central to our considerations because the atom will effectively behave as a clock at a superposition of heights within the optical lattice.
-The presence of two clock pulses is the main modification to previous experimental implementations [18, 19, 20, 21, 22].
-Let us notice, that without those pulses, the probability of measuring the atoms in the ground state would be constant and equal to 1/4 (recall that we lose some of the trajectories).
-In our setup, they oscillate between 0 and 1/4, depending on the time $T_{\text{B}}$ of the lattice hold.
+Apart from the non-relativistic phase $\Delta\phi$ between the upper and lower ground-state trajectories (measured in typical interferometric experiments, e.g.
+[20]), there will be additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired within the optical lattice by the excited state on the lower and upper trajectories and which can be measured with the proposed setup.
+The probability $P_{g}^{(j)}$ is given by (see Appendix A for derivation):
+$P_{g}^{(j)}=\frac{1}{16}\bigg{[}1-\cos\left(\frac{\delta_{\rm u}+\delta_{\rm d}}{2}\right)\cos\left(\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)+2(-1)^{j}\cos\left(\Delta\phi+\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm u}}{2}\right)\bigg{]}.$ (1)
+To eliminate the $\Delta\phi$-dependence, we calculate the total probability $P_{g}=P_{g}^{(0)}+P_{g}^{(1)}$ of finding the atom in the ground state
+$P_{g}=\frac{1}{8}\left[1-\cos\left(\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\cos\left(\frac{\delta_{\rm u}+\delta_{\rm d}}{2}\right)\right].$ (2)
+This probability depends only on the evolution confined to the optical-lattice region, where the atom effectively behaves as a clock in a superposition of heights.
+The presence of two clock pulses is the sole modification to previous implementations [18, 19, 20, 21, 22].
+Note that without these pulses, the ground-state probability would remain fixed at 1/4 (recall that we lose some of the trajectories).
+In our setup, they oscillate between 0 and 1/4 as a function of the time $T_{\rm B}$ between the clock pulses.
 —Dynamics and phases.
-Let us analyze the optical lattice stage of the interferometer and the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state.
-At this stage, the atom will Bloch oscillate at a superposition of heights separated by $\Delta z$.
+We now analyze the optical-lattice stage of the interferometer and the additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired by the excited state.
+During this stage, the atom undergoes Bloch oscillations in a superposition of two heights separated by $\Delta z$.
 We adopt a simplified model of Bloch oscillations (see [61, 19] and Fig.
-2), in which the atom interacts with the optical lattice only when it reaches a downward velocity $v_{\text{B}}=\hbar k/m$ (here $k$ is the wave vector of the optical-lattice laser and $m$ is the mass of the ground-state atom).
-Then, it instantaneously exchanges two photons with the lattice and changes its velocity.
-For the appropriately chosen frequency of the lattice light, the velocity at the interaction point will simply flip its sign, and the atom will oscillate at the fixed height with period $\tau_{\text{B}}=2v_{\text{B}}/g$.
+2), in which the atom interacts with the optical lattice only when its downward velocity reaches $v_{\rm B}=\hbar k/m$ (here $k$ is the wave vector of the optical-lattice laser and $m$ is the mass of the ground-state atom).
+At that point, the atom instantaneously exchanges two photons with the lattice.
+For an appropriately chosen lattice-light frequency, this leads to velocity reversal occurring always at the same height, causing the atom to oscillate with period $\tau_{\rm B}=2v_{\rm B}/g$.
 We will assume that the laser frequency has been chosen exactly this way.
-Each bounce starts and ends at the apex of a free-fall trajectory.
-Within the time $T_{\text{B}}$ the atom performs $N$ oscillations with the last one possibly extended (or shortened) by a time $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$ (this is because $T_{\text{B}}$ is not necessarily an integer multiple of the Bloch period $\tau_{\text{B}}$; see Fig.
-2).
-$N\tau_{\text{B}}$$\tau$$T_{\text{B}}$
+Each oscillation thus begins and ends at the apex of a free-fall trajectory.
+Over the hold time $T_{\rm B}$, the atom completes $N$ full oscillations, with the final one possibly extended or shortened by an additional time $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$, reflecting that $T_{\rm B}$ need not be an integer multiple of $\tau_{\rm B}$.
+[…]
 Figure 2: Trajectory followed by the atom (at the lower or upper trajectory) in the optical lattice.
 We adopt a simplified model of Bloch oscillations in which the atom follows a bouncing trajectory at a fixed height.
-The total time of Bloch oscillations $T_{\text{B}}$ needs not be equal to an integer multiple of Bloch periods $N\tau_{\text{B}}$.
-Instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\pi/2,\pi/2)$.
-Crucially, due to the mass difference between the excited and the ground state of the atom, the condition for bouncing at the fixed height can be satisfied only for one of the internal states for a given lattice.
-For instance, if we choose the lattice laser frequency so that the condition is satisfied for the ground state, the excited state will necessarily fall beyond the starting height with each oscillation.
-However, it can be shown (see Appendix B) that in viable experimental scenarios, the total fall is much smaller than the thermal wavelength of the atom.
-Therefore, employing the perturbative approach developed in [58] is justified, which here means that we can approximate that both the ground and the excited state follow the same trajectory.
-We further verify that this approximation holds in Appendix D.
-The phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ consist of two components: laser phase $\omega T_{\text{B}}$ acquired by the excited state due to the clock pulses, and a propagation phase correction due to different masses of the ground and the excited state.
-To calculate the latter we follow [58] and express the Lagrangian of the excited state $L_{e}=L_{g}+\delta L$ in terms of the Lagrangian $L_{g}$ of the ground state and a small correction $\delta L$ resulting from the rest-mass difference between the ground and the excited state.
-If we denote this mass difference by $\delta m=\hbar\omega_{0}/c^{2}$, the Lagrangian correction is given by
+The total time of Bloch oscillations $T_{\rm B}$ need not be equal to an integer multiple of Bloch periods $N\tau_{\rm B}$.
+Instead, $T_{\rm B}=N\tau_{\rm B}+\tau$ with $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$.
+Crucially, due to the mass difference between the excited and ground states of the atom, the condition for oscillations at the fixed height can be satisfied only for one internal state at a time for a given lattice frequency.
+If the lattice is tuned to satisfy this condition for the ground state, the excited state will fall below the starting height after each oscillation.
+However, as shown in Appendix B, in viable experimental scenarios, the cumulative fall distance is much smaller than the atom’s thermal de Broglie wavelength.
+Therefore, the perturbative approach of [62] is justified, allowing us to approximate that both internal states follow the same trajectory.
+This approximation is further validated in Appendix D.
+The phases $\delta_{\rm d}$ and $\delta_{\rm u}$ comprise two contributions: the laser phase $\omega T_{\rm B}$ acquired by the excited state due to the clock pulses, and a propagation-phase correction arising from the different masses of the ground and excited states.
+To calculate the latter, we follow [62] and write the excited-state Lagrangian as $L_{e}=L_{g}+\delta L$, where $L_{g}$ is the ground-state Lagrangian and $\delta L$ is a small correction resulting from the rest-mass difference between the two states.
+If we denote this mass difference by $\delta m=\hbar\omega_{0}/c^{2}$, the Lagrangian correction is given by:
 $\delta L=-\delta m\left(c^{2}+gz-v^{2}/2\right).$ (3)
-Then, to calculate the propagation phase correction we integrate $\delta L$ over the trajectory followed by the atom and divide by $\hbar$.
-The full calculation is presented in the Appendix C (followed by a non-perturbative calculation in Appendix D).
-It leads to the following formulae for the phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$:
-$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}},\end{split}$ (4)
-where $\langle v^{2}\rangle=v_{\text{B}}^{2}/3$ is the mean square speed within a single oscillation.
-In principle, the mass-energy corrections to the interferometric phases are not the only corrections of the order of $1/c^{2}$ that are present in the proposed experiment — following [59] we could also include corrections coming from the higher-order expansion of the free-fall Lagrangian of the relativistic particle, and $1/c^{2}$ corrections to the laser phases coming from elastic interactions (Bragg and Bloch pulses).
-These depend only on the trajectory followed by the atom and thus, in the picture in which ground and excited states follow the same trajectories, affect only the interferometric phase $\Delta\phi$ leaving $\delta_{\text{k}}$ and $\delta_{\text{g}}$ unmodified.
-However, the $1/c^{2}$ corrections to $\Delta\phi$ are not the main focus of our considerations and we will not deal with them.
-On the contrary, any corrections to the clock-pulse laser phases would affect $\delta_{\text{k}}$ and $\delta_{\text{g}}$ if they were present.
-Let us notice, however, that the corrections to the laser phases calculated in [59] resulted from the assumption that the atom-light interactions are always resonant and the atom interacts with photons with laboratory-frame frequencies depending on its position and velocity.
-For our present purposes, we assume instead that the laser frequency of the clock pulses is fixed and the non-resonant processes are possible.
-Then, there are no corrections to the clock-pulse laser phases because the atom interacts with photons with the same laboratory-frame frequency regardless of its height and velocity.
+To calculate the propagation-phase correction, we integrate $\delta L$ along the atom’s trajectory and divide by $\hbar$.
+The full derivation is given in Appendix C; a non-perturbative calculation is presented in Appendix D.
+This yields the expressions for the phases $\delta_{\rm d}$ and $\delta_{\rm u}$:
+$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\rm B},\end{split}$ (4)
+where $\langle v^{2}\rangle=v_{\rm B}^{2}/3$ is the mean square speed within a single oscillation.
+In principle, mass–energy corrections to the interferometric phases are not the only $\mathcal{O}(1/c^{2})$ effects present in our proposed experiment.
+Following Werner et al.
+[63], one could also include higher-order corrections from the expansion of the free-fall Lagrangian of the relativistic particle, and $\mathcal{O}(1/c^{2})$ corrections to the laser phases arising from elastic interactions (Bragg and Bloch pulses).
+Because these contributions depend only on the atomic trajectory, and since we approximate that both internal states follow the same path, they modify only the interferometric phase $\Delta\phi$, leaving $\delta_{\rm d}$ and $\delta_{\rm u}$ unmodified.
+However, $\mathcal{O}(1/c^{2})$ corrections to $\Delta\phi$ lie beyond the main focus of this work and will be neglected.
+Conversely, any corrections to the clock-pulse phases would affect $\delta_{\rm d}$ and $\delta_{\rm u}$ if present.
+Note that the laser-phase corrections calculated in [63] arise from assuming resonant atom-light interactions, with the atom interacting with photons whose laboratory-frame frequencies depend on its position and velocity.
+Here, we instead assume the clock pulses operate at a fixed laboratory-frame frequency, permitting non-resonant processes.
+Under this assumption, there are no corrections to the clock-pulse phases, since the atom interacts with photons at the same frequency regardless of its height or velocity.
 —Results and discussion.
-Let us denote $\varepsilon_{\text{k}}=\langle v^{2}\rangle/2c^{2}$ for the kinetic correction in Eq.
-(4) and $\varepsilon_{\text{g}}=g\Delta z/c^{2}$ for the gravitational correction, and denote by $\langle\omega_{0}\rangle=\omega_{0}\left(1+\varepsilon_{\text{k}}+\varepsilon_{\text{g}}/2\right)$ the (quantum mechanical) mean of the atomic internal frequency.
-Then, the probability of detecting the atom in the ground state (compare with Eq.
-(2)) can be written as
-$P_{g}=\frac{1}{8}\left[1-\mathcal{V}(T_{\text{B}})\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)\right],$ (5)
-where $\mathcal{V}(T_{\text{B}}):=\cos\left(\varepsilon_{\text{g}}\omega_{0}T_{\text{B}}/2\right)$ is the interferometric visibility.
-In a generic situation (away from the resonance), the visibility varies very slowly with time $T_{\text{B}}$ compared to $\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)$.
-Therefore, Eq.
-(5) describes an interference pattern with a slowly varying envelope.
+Let us introduce $\varepsilon_{\rm k}=\langle v^{2}\rangle/2c^{2}$ for the kinetic correction in Eq.
+(4) and $\varepsilon_{\rm g}=g\Delta z/c^{2}$ for the gravitational correction.
+Denoting the quantum-mechanical mean internal frequency by $\langle\omega_{0}\rangle=\omega_{0}\left(1-\varepsilon_{\rm k}+\varepsilon_{\rm g}/2\right)$, the ground-state detection probability (cf.
+Eq.
+(2)) can be written as:
+$P_{g}=\frac{1}{8}\left[1-\mathcal{V}(T_{\rm B})\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\rm B}\right)\right],$ (5)
+where $\mathcal{V}(T_{\rm B}):=\cos\left(\varepsilon_{\rm g}\omega_{0}T_{\rm B}/2\right)$ is the interferometric visibility.
+In a generic operating regime away from resonance, the visibility envelope varies slowly with time $T_{\rm B}$ compared to the rapid oscillations of $\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\rm B}\right)$.
+Thus, Eq.
+(5) yields an interference pattern comprising fast oscillations modulated by a slowly varying envelope.
 Apart from the visibility modulations, Eq.
 (5) looks like a result of a Ramsey interferometry experiment for an atom with internal transition frequency $\langle\omega_{0}\rangle$.
-For fixed time $T_{\text{B}}$ one can scan through different clock pulse frequencies $\omega$ to find this resonant frequency.
-Thus, the relativistic effects coming from mass-energy equivalence can be measured in two ways: by observing the visibility oscillations, and by measuring the fractional frequency shift
-$\left\langle\frac{\delta\omega_{0}}{\omega_{0}}\right\rangle\equiv\frac{\langle\omega_{0}\rangle-\omega_{0}}{\omega_{0}}=\varepsilon_{\text{k}}+\frac{\varepsilon_{\text{g}}}{2}.$ (6)
-Note that while the visibility oscillations are sensitive only to the gravitational correction $\varepsilon_{\text{g}}$, the frequency shift depends both on $\varepsilon_{\text{g}}$ and $\varepsilon_{\text{k}}$.
-To distinguish between the gravitational and the kinetic effect in fractional frequency shift measurements, one can modulate the separation $\Delta z$ between the upper and the lower trajectory — this will affect the gravitational correction, but not the kinetic one.
-The probability $P_{g}$ as a function of time is visualized in Fig.
-3 for exemplary values of $\omega$, $\omega_{0}$, $\varepsilon_{\text{g}}$, and $\varepsilon_{\text{k}}$.
-[…]
-Figure 3: Oscillations of the (normalized) probability $P_{g}$ for exemplary values $\omega=40\pi$, $\omega_{0}=1.1\omega$, and $\varepsilon_{\text{k}}=0.01$, and two different values of the gravitational correction: $\varepsilon_{\text{g}}=0$ (in gray) and $\varepsilon_{\text{g}}=\varepsilon_{\text{k}}$ (in orange).
+For a fixed $T_{\rm B}$, scanning the clock-pulse frequency $\omega$ reveals the resonance at $\omega=\langle\omega_{0}\rangle$.
+Consequently, relativistic effects arising from mass–energy equivalence can be probed in two ways: by observing visibility oscillations, and by measuring the resulting fractional frequency shift
+$\left\langle\frac{\delta\omega_{0}}{\omega_{0}}\right\rangle\equiv\frac{\langle\omega_{0}\rangle-\omega_{0}}{\omega_{0}}=-\varepsilon_{\rm k}+\frac{\varepsilon_{\rm g}}{2}.$ (6)
+Note that whereas visibility oscillations are sensitive only to the gravitational correction $\varepsilon_{\rm g}$, the frequency shift depends on both $\varepsilon_{\rm g}$ and $\varepsilon_{\rm k}$.
+To distinguish the gravitational and kinetic contributions in fractional-frequency-shift measurements, one can modulate the height separation $\Delta z$, which alters $\varepsilon_{\rm g}$ but not $\varepsilon_{\rm k}$.
+The probability $P_{g}$ as a function of time is shown in Fig.
+3 for exemplary values of $\omega$, $\omega_{0}$, $\varepsilon_{\rm g}$, and $\varepsilon_{\rm k}$.
+Assuming the experimental parameters as in [20], we would have the velocity $v_{\rm B}\sim 10\,\text{mm/s}$, and the vertical separation of the trajectories $\Delta z\sim 10\,\mu\text{m}$.
+This would mean that the corrections would be of the order $|\varepsilon_{\rm k}|\sim|\varepsilon_{\rm g}|\sim 10^{-22}$.
+The fractional frequency shift of that order is very close to what can be currently detected [64, 52, 53, 65].
+Indeed, state-of-the-art atomic clocks reach the multiple-measurement precision of the order of $10^{-21}$ [52, 53, 65] (the lowest single-measurement precision in such experiments is $\sim 10^{-19}$ [66]), which is the order of magnitude for the fractional frequency shift in the proposed experiment for separations $\Delta z\sim 100\,\mu\text{m}$, which has already been achieved in [20], at the expense of shorter coherence time.
+Figure 3: Oscillations of the (normalized) probability $P_{g}$ for exemplary values $\omega=40\pi$, $\omega_{0}=1.1\omega$, and $\varepsilon_{\rm k}=0.01$, and two different values of the gravitational correction: $\varepsilon_{\rm g}=0$ (in gray) and $\varepsilon_{\rm g}=\varepsilon_{\rm k}$ (in orange).
 Note that, apart from visibility modulations, the orange pattern has its frequency shifted against the gray one.
-[…]
-Note that the observation of the frequency shift (6) itself would not be evidence of the quantum clock reading the superposition of proper times since the frequency shift is the same as for the clock being in a statistical mixture of two heights.
-To reject the latter possibility one can, apart from the total probability $P_{g}$ of measuring the atom in the ground state at the output, extract from the measured data also the difference $D_{g}\equiv P_{g}^{(0)}-P_{g}^{(1)}$ given by (compare with Eq.
+On the other hand, for a transition frequency $\omega_{0}\sim 10^{15}\,\mathrm{Hz}$ and a lattice-hold time $T_{\rm B}\sim 1\,\mathrm{s}$ (as in [20]), with $|\varepsilon_{\rm g}|\sim 10^{-22}$, the predicted visibility drop is of order $10^{-15}$.
+This is far below current experimental sensitivity.
+Although strontium atoms can support coherent Bloch oscillations for over $100\,\mathrm{s}$ [67], the longest achieved to date is $\sim 1\,\mathrm{s}$.
+Minute-scale oscillation times have been demonstrated with cesium atoms [21, 22], but their lower $\omega_{0}$ renders the effect even weaker.
+Extending $T_{\rm B}$ to $100\,\mathrm{s}$ would amplify the visibility oscillations by $10^{4}$, yet they would remain negligible.
+To achieve an observable visibility drop (of order 0.0110.01-1), one would need a trajectory separation of $\Delta z\sim 10-100\,\mathrm{cm}$.
+Such separations exceed current capabilities and would likely require two independent lattice lasers.
+Note that the observation of the frequency shift (6) itself would not be evidence of the quantum clock reading the superposition of proper times, since the frequency shift is the same as for the clock being in a statistical mixture of two heights.
+To reject the latter possibility, one can, apart from the total probability $P_{g}$ of measuring the atom in the ground state at the output, extract from the measured data also the difference $D_{g}\equiv P_{g}^{(0)}-P_{g}^{(1)}$ given by (compare with Eq.
 (1))
-$D_{g}=\frac{1}{8}\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right),$ (7)
-where $\delta_{\text{d}}$ and $\delta_{\text{u}}$ are given in Eq. (4), and $\Delta\phi$ is given by (see Appendix C for derivation)
+$D_{g}=\frac{1}{8}\cos\left(\Delta\phi+\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm u}}{2}\right),$ (7)
+where $\delta_{\rm d}$ and $\delta_{\rm u}$ are given in Eq. (4), and $\Delta\phi$ reads (see Appendix C for derivation)
 $\Delta\phi=-\frac{mg\Delta z}{\hbar}\left(T+2T^{\prime}+\tau\right).$ (8)
-If the atom is in a superposition of heights, $D_{g}$ will oscillate with time $T_{\text{B}}$.
+If the atom is in a superposition of heights, $D_{g}$ will oscillate with time $T_{\rm B}$.
 On the other hand, for the atom following a statistical mixture of two heights, the probability difference would be equal to zero.
 Therefore, extracting from the interferometric output the fractional frequency shift and the oscillations of $D_{g}$ would provide evidence that the atom was reading a superposition of proper times.
-It is important to emphasize that the method proposed in this work does not rely on the assumption that ground and excited states follow the same trajectories within the optical lattice.
-While we showed that within the required order $(1/c^{2})$ of calculating the phases acquired in the interferometer the trajectories associated with different mass-energy states can be approximated as equal, the method allows accounting for mass-energy corrections and resulting phases beyond this approximation.
-Finally, let us notice that the interpretation of the experiment as a spectroscopy of a clock at the superposition of heights is possible thanks to including the laser phases and the concrete mechanism that sets up the quantum clock.
-These were missing in the idealized scenarios in [23, 24] and in the recent work [38] proposing a similar setup using optical tweezers.
-Therefore, the possibility of measuring the relativistic frequency shift in such interferometric scenarios thus far went unnoticed.
-Acknowledgements.
-We thank Timothy Le Pers, Navdeep Arya, Germain Tobar, and Jun Ye for helpful discussions and useful comments.
-The authors acknowledge Knut and Alice Wallenberg foundation through a Wallenberg Academy Fellowship No.
+It is important to emphasize that the method proposed in this work does not rely on the assumption that the ground and excited states follow the same trajectories within the optical lattice.
+While we showed that, within the required order $1/c^{2}$ of calculating the phases acquired in the interferometer, the trajectories associated with different mass–energy states can be approximated as equal, the method allows accounting for mass–energy corrections and the resulting phases beyond this approximation.
+Furthermore, we emphasize that our results do not depend crucially on the simplified Bloch oscillation model we have adopted.
+Indeed, in Appendix E, we derive analogous results using the usual description of Bloch oscillations [68, 69], in which the atom oscillates under a periodic potential of the optical lattice.
+The assumption of zero momentum transfer from the clock pulse is well motivated in special cases, including the usual Lamb-Dicke regime of optical-lattice traps [54, 56, 57, 58], and is exact in the case of two-photon E1-M1 clock transitions [55, 59, 60].
+The Lamb-Dicke regime is reached when the atomic motion is limited to a region smaller than the transition wavelength.
+It was shown [54] that in this case, most transitions occur at the internal transition frequency and do not alter the center-of-mass motion of the atom.
+However, such behavior is well understood only in the fully quantum picture (with quantized center-of-mass motion) — in our simplified picture with the atom following a semiclassical trajectory, the no-recoil assumption is made as a simplifying yet justified approximation.
+By contrast, for E1-M1 transitions [55, 59, 60], involving two counter-propagating photons with equal frequencies, the clock transition is recoilless as a rule.
+In this case, the no-recoil assumption is also well motivated in our simplified model of Bloch oscillations.
+Abandoning the assumption of zero momentum transfer naturally leads to recoil shifts [70, 71].
+In our model, this implies that the ground and excited states follow different trajectories, and the description becomes significantly more involved.
+The analysis of this case goes beyond the scope of the present work.
+Note that the interpretation of the experiment as spectroscopy of a clock in a superposition of heights is possible only by including the laser phases and the explicit mechanism that establishes the quantum clock.
+These elements were absent in the idealized scenarios of [23, 24] and in the recent proposal using optical tweezers [38].
+Consequently, the possibility of measuring the relativistic frequency shift in such interferometric setups has gone unnoticed until now.
+We thank Timothy Le Pers, Navdeep Arya, Germain Tobar, Jun Ye, Kyungtae Kim, and Charles Baynham for helpful discussions and useful comments.
+The authors acknowledge the Knut and Alice Wallenberg Foundation through a Wallenberg Academy Fellowship No.
 2021.0119.
@@
 Appendix A Probability calculation
-Here we derive the formula (1) for the probability $P_{g}^{(j)}$ of measuring the atom in the ground state in output $j\in\{0,1\}$.
-We denote by $\Delta\phi$ the phase difference between upper and lower ground state trajectories, and by $\delta_{\text{d}}$ and $\delta_{\text{u}}$ the additional phase acquired by the excited state on the lower and upper trajectory, respectively.
-The final state of the atom |$\ket{\psi_{\text{final}}}$ can be written in a generic way
-|$\ket{\psi_{\text{final}}}=\frac{1}{8}\ket{g}\left(\ket{0}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]-\mathrm{i}\ket{1}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}-\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]\right)+\ldots,$ (9)
-where |$\ket{0}$ and |$\ket{1}$ are the states corresponding to two relevant output trajectories, and dots at the end stand for the parts of the state irrelevant to our considerations (lost trajectories and excited state).
-The factor $1/8=(1/\sqrt{2})^{6}$ in front comes from six laser pulses applied to the atom (four Bragg pulses and two clock pulses).
-The minus signs in the bracket multiplying |$\ket{0}$ result from the fact that excited state trajectories interacted with a laser two more times than the ground state trajectories (two clock pulses) — each interaction gives rise to the factor $\mathrm{i}\mathrm{e}^{\mathrm{i}\phi_{l}}$ multiplying the corresponding trajectory (here $\phi_{l}$ is the so-called laser phase and we absorb it into phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$).
-The factor $-\mathrm{i}$ multiplying |$\ket{1}$ reflects the fact that trajectories leaving the interferometer in output 1 interacted with the odd number of Bragg pulses (the lower trajectory interacted once, the upper one three times), while the ones leaving in output 0 interacted with the even number (two) of Bragg pulses.
-Finally, the additional minus sign in the bracket multiplying |$\ket{1}$ comes from the fact that, in the case of trajectories leaving in output 1, the upper trajectory interacted with two more Bragg pulses than the lower one.
+[…]
+[…]
+[…]
 The amplitude of finding the atom in the ground state in the output $j$ is given by
-$\bra{g,j}\ket{\psi_{\text{final}}}\propto\frac{1}{8}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+(-1)^{j}\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right],$ (10)
-where we omitted the common phase factor $-\mathrm{i}$ present in the case of $j=1$. To calculate the probability $P_{g}^{(j)}$ we take the absolute value squared of the above amplitude and get
-$P_{g}^{(j)}=\left|\bra{g,j}\ket{\psi_{\text{final}}}\right|^{2}=\frac{1}{16}\left[1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\right],$ (11)
+[…]
+where we omitted the common phase factor $\mathrm{i}$ present in the case of $j=1$. To calculate the probability $P_{g}^{(j)}$ we take the absolute value squared of the above amplitude and get
+[…]
 which is exactly Eq. (1).
@@
 Let us motivate the use of the perturbative method [62] by analyzing the fall of the excited-state trajectory during Bloch oscillations.
-We assume that the frequencies of the optical lattice lasers are chosen such that the ground state follows a bouncing trajectory at a fixed height (i.e., each interaction with the laser occurs at the same height $z_{\text{B}}$).
+We assume that the frequencies of the optical lattice lasers are chosen such that the ground state follows a bouncing trajectory at a fixed height (i.e., each interaction with the laser occurs at the same height $z_{\rm B}$).
 Then, the excited-state trajectory necessarily falls with each bounce, due to its greater mass.
-Indeed, since we assume that the interaction with the laser occurs only when the atom reaches downward velocity $v_{\text{B}}=\hbar k/m$, and the momentum transferred to the atom in each interaction is the same for both the ground and the excited state, the excited state will move after each interaction with the upward velocity $v_{\text{B}}-\delta v_{\text{B}}$, where the velocity difference $\delta v_{\text{B}}$ is given by
-$\delta v_{\text{B}}=2v_{\text{B}}\frac{\delta m}{m}.$ (12)
+Indeed, since we assume that the interaction with the laser occurs only when the atom reaches downward velocity $v_{\rm B}=\hbar k/m$, and the momentum transferred to the atom in each interaction is the same for both the ground and the excited state, the excited state will move after each interaction with the upward velocity $v_{\rm B}-\delta v_{\rm B}$, where the velocity difference $\delta v_{\rm B}$ is given by
+$\delta v_{\rm B}=2v_{\rm B}\frac{\delta m}{m}.$ (12)
 Since the velocity of the excited state right before each interaction is larger (in terms of the absolute value) than right after it, each subsequent interaction will occur lower by
-$\delta z_{\text{B}}=\frac{v_{\text{B}}\delta v_{\text{B}}}{g}$ (13)
+$\delta z_{\rm B}=\frac{v_{\rm B}\delta v_{\rm B}}{g}$ (13)
 than the previous one meaning that the excited-state trajectory will progressively fall.
-Also, the excited trajectory’s smaller starting velocity means it will reach the downward velocity $v_{\text{B}}$ faster than the ground state.
-Two subsequent interactions on the excited-state trajectory are separated by a time $\tau_{\text{B}}-\delta\tau_{\text{B}}$, where
-$\delta\tau_{\text{B}}=\frac{\delta v_{\text{B}}}{g}.$ (14)
+Also, the excited trajectory’s smaller starting velocity means it will reach the downward velocity $v_{\rm B}$ faster than the ground state.
+Two subsequent interactions on the excited-state trajectory are separated by a time $\tau_{\rm B}-\delta\tau_{\rm B}$, where
+$\delta\tau_{\rm B}=\frac{\delta v_{\rm B}}{g}.$ (14)
 The spatial separation between the ground and the excited state trajectories is maximal when the excited state interacts with the laser.
-Then, it changes the direction of motion and meets the ground state trajectory at its interaction height $z_{\text{B}}$.
-Therefore, the trajectories meet repeatedly (at the interaction points of the ground state trajectory), but at each subsequent meeting point the excited-state trajectory has velocity lower by $\delta v_{\text{B}}$ than at the previous meeting point (while the ground state moves there at velocity $v_{\text{B}}$).
-[…]
+Then, it changes the direction of motion and meets the ground state trajectory at its interaction height $z_{\rm B}$.
+Therefore, the trajectories meet repeatedly (at the interaction points of the ground state trajectory), but at each subsequent meeting point the excited-state trajectory has velocity lower by $\delta v_{\rm B}$ than at the previous meeting point (while the ground state moves there at velocity $v_{\rm B}$).
+The largest separation between the $n$th and $(n+1)$th interaction on the ground state trajectory occurs at the moment of $(n+1)$th interaction on the excited-state trajectory and is equal to $2n\delta z_{\rm B}$.
+For the experimental parameters from [20] (strontium atoms, optical lattice operating at $532\;\text{nm}$, around 500 oscillations) the maximal separation between the ground and excited state trajectories would be around $10^{-13}\;\text{m}$.
+This should be compared with the thermal wavelength of the atoms given by
 $\lambda_{\text{th}}=\sqrt{\frac{2\pi\hbar^{2}}{mk_{\text{\tiny{B}}}T}},$ (15)
@@
 Appendix C Phases
-Let us calculate the phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$ using the perturbative method developed in [58].
+Let us calculate the phases $\Delta\phi$, $\delta_{\rm d}$, and $\delta_{\rm u}$ using the perturbative method developed in [62].
 Each consists of two contributions: the propagation phase calculated by integrating the free-fall Lagrangian over the trajectory followed by the atom, and the laser phase resulting from atom-light interactions.
@@
 Note also that the correction $\delta L^{(2)}$ and the corrections to the laser phases derived in [63] depend only on the trajectory, not on the internal state.
-Therefore, in the picture in which both the ground and the excited state follow the same trajectory, they will contribute only to the phase $\Delta\phi$, not to $\delta_{\text{d}}$ and $\delta_{\text{u}}$.
+Therefore, in the picture in which both the ground and the excited state follow the same trajectory, they will contribute only to the phase $\Delta\phi$, not to $\delta_{\rm d}$ and $\delta_{\rm u}$.
 We are not interested in $1/c^{2}$ corrections to $\Delta\phi$, hence we will omit them in the subsequent analysis and focus only on the corrections coming from $\delta L^{(1)}$.
@@
 $L_{g}=-m\left(c^{2}+gz-v^{2}/2\right)$ (19)
-along the trajectory. More precisely, for the free fall starting at time $t_{\text{A}}$ and ending at $t_{\text{B}}$, the propagation phase is given by
-$\begin{split}\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=&\frac{1}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}tL_{g}=-\frac{m}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}t\left(c^{2}+gz-v^{2}/2\right)\\ =&-\frac{m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right],\end{split}$ (20)
+along the trajectory. More precisely, for the free fall starting at time $t_{\text{A}}$ and ending at $t_{\rm B}$, the propagation phase is given by
+$\begin{split}\phi_{\text{p}}(t_{\text{A}}\to t_{\rm B})=&\frac{1}{\hbar}\int_{t_{\text{A}}}^{t_{\rm B}}\mathrm{d}tL_{g}=-\frac{m}{\hbar}\int_{t_{\text{A}}}^{t_{\rm B}}\mathrm{d}t\left(c^{2}+gz-v^{2}/2\right)\\ =&-\frac{m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\rm B}-t_{\text{A}})+v_{\text{A}}g(t_{\rm B}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\rm B}-t_{\text{A}})^{3}\right],\end{split}$ (20)
 where $z_{\text{A}}$ and $v_{\text{A}}$ are the atom’s position and velocity, respectively, at time $t_{\text{A}}$.
@@
 Notice that to match the notation from the main text, the times of application of particular pulses must satisfy
-$t_{2}-t_{1}=t_{4}-t_{3}=T,\qquad t_{\text{i}}-t_{2}=t_{3}-t_{\text{f}}=T^{\prime},\qquad t_{\text{f}}-t_{\text{i}}=T_{\text{B}}.$ (21)
-Let us further denote the quantities corresponding to the lower and upper trajectories by superscripts (d) and (u), respectively, and by $\Delta\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=\phi_{\text{p}}^{\text{(u)}}(t_{\text{A}}\to t_{\text{B}})-\phi_{\text{p}}^{\text{(d)}}(t_{\text{A}}\to t_{\text{B}})$ the propagation phase difference between those two trajectories acquired between $t_{\text{A}}$ and $t_{\text{B}}$.
+$t_{2}-t_{1}=t_{4}-t_{3}=T,\qquad t_{\text{i}}-t_{2}=t_{3}-t_{\text{f}}=T^{\prime},\qquad t_{\text{f}}-t_{\text{i}}=T_{\rm B}.$ (21)
+Let us further denote the quantities corresponding to the lower and upper trajectories by superscripts (d) and (u), respectively, and by $\Delta\phi_{\text{p}}(t_{\text{A}}\to t_{\rm B})=\phi_{\text{p}}^{\text{(u)}}(t_{\text{A}}\to t_{\rm B})-\phi_{\text{p}}^{\text{(d)}}(t_{\text{A}}\to t_{\rm B})$ the propagation phase difference between those two trajectories acquired between $t_{\text{A}}$ and $t_{\rm B}$.
 Let us consider one by one the consecutive stages of the interferometer and calculate the corresponding phase difference using (20).
@@
 The relative propagation phase equals
-    $\Delta\phi_{\text{p}}(t_{\text{i}}\to t_{\text{f}})=-\frac{m}{\hbar}g\Delta vTT_{\text{B}}.$ (24)
+    $\Delta\phi_{\text{p}}(t_{\text{i}}\to t_{\text{f}})=-\frac{m}{\hbar}g\Delta vTT_{\rm B}.$ (24)
   • $t_{\text{f}}\to t_{3}$:
@@
 To calculate the total propagation phase difference $\Delta\phi_{\text{p}}$ we sum up all the contributions, which gives
-$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta vT\left[v_{1}+\Delta v-v_{3}-g\left(T+2T^{\prime}+T_{\text{B}}\right)\right].$ (27)
+$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta vT\left[v_{1}+\Delta v-v_{3}-g\left(T+2T^{\prime}+T_{\rm B}\right)\right].$ (27)
 Finally, let us note that, because of the requirement that at $t_{\text{i}}$ the velocity on both trajectories is equal to zero, the following relation holds
@@
 Let us notice that the atom does not leave the optical lattice with zero velocity (see Fig.
-2), but rather with velocity $-g\tau$.
-Here $\tau\in\{-\tau_{\text{B}}/2,\tau_{\text{B}}/2\}$ is given by
-$T_{\text{B}}=N\tau_{\text{B}}+\tau,$ (29)
-where $N=\lfloor T_{\text{B}}/\tau_{\text{B}}+1/2\rfloor$ is the number of interactions between the optical lattice and the atom. Therefore, we have
+2), but rather with velocity $g\tau$.
+Here $\tau\in\{-\tau_{\rm B}/2,\tau_{\rm B}/2\}$ is given by
+$T_{\rm B}=N\tau_{\rm B}+\tau,$ (29)
+where $N=\lfloor T_{\rm B}/\tau_{\rm B}+1/2\rfloor$ is the number of interactions between the optical lattice and the atom. Therefore, we have
 $v_{3}=-g(\tau+T^{\prime})$ (30)
 and we can rewrite Eq. (27) (introducing $\Delta z=\Delta vT$) as follows:
-$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta zg(\tau-T_{\text{B}})=-\frac{m}{\hbar}g\Delta zN\tau_{\text{B}}.$ (31)
+$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta zg(\tau-T_{\rm B})=-\frac{m}{\hbar}g\Delta zN\tau_{\rm B}.$ (31)
 Now, let us focus on the laser phase contribution to the phase $\Delta\phi$.
 As described in [62], when the atom absorbs (emits) a photon with frequency $\omega$ and wave vector $k$, it acquires a laser phase
-$\phi_{\text{l}}(t,z)=\mp\mathrm{i}\omega t\pm kz,$ (32)
+$\phi_{\text{l}}(t,z)=\mp\omega t\pm kz,$ (32)
 where the upper (lower) sign corresponds to photon absorption (emission), and $t$ and $z$ are the time and position of the interaction, respectively.
@@
 Therefore, the laser phase acquired at each such interaction point is equal to $\phi_{\text{l}}(t,z)=\phi_{\text{l}}(z)=\pm 2kz$ with the plus (minus) sign corresponding to the atom receiving upward (downward) momentum.
-Denote by $z_{j}$ with $j\in\{1,2,3,4\}$ the interaction heights with four consecutive Bragg pulses ($z_{1}$ and $z_{3}$ correspond to the upper trajectory, while $z_{2}$ and $z_{4}$ to the lower one), and by $z_{\text{d}}$ and $z_{\text{u}}$ the heights of interactions of the lower and upper trajectories, respectively, within the optical lattice.
+Denote by $z_{j}$ with $j\in\{1,2,3,4\}$ the interaction heights with four consecutive Bragg pulses ($z_{1}$ and $z_{3}$ correspond to the upper trajectory, while $z_{2}$ and $z_{4}$ to the lower one), and by $z_{\rm d}$ and $z_{\rm u}$ the heights of interactions of the lower and upper trajectories, respectively, within the optical lattice.
 Note also that the wave vector $k_{\text{Bragg}}$ of the pulses is related to the velocity change $\Delta v$ by
 $2\hbar k_{\text{Bragg}}=m\Delta v,\qquad\Longrightarrow\qquad k_{\text{Bragg}}=\frac{m\Delta v}{2\hbar},$ (33)
-and the wave vector $k_{\text{Bloch}}$ of the optical lattice laser is related to the Bloch period $\tau_{\text{B}}$ by
-$2\hbar k_{\text{Bloch}}=mg\tau_{\text{B}},\qquad\Longrightarrow\qquad k_{\text{Bloch}}=\frac{mg\tau_{\text{B}}}{2\hbar}.$ (34)
+and the wave vector $k_{\text{Bloch}}$ of the optical lattice laser is related to the Bloch period $\tau_{\rm B}$ by
+$2\hbar k_{\text{Bloch}}=mg\tau_{\rm B},\qquad\Longrightarrow\qquad k_{\text{Bloch}}=\frac{mg\tau_{\rm B}}{2\hbar}.$ (34)
 The total laser phase difference (between the upper and lower trajectory) is given by
-$\Delta\phi_{\text{l}}=2k_{\text{Bragg}}(z_{1}-z_{2}-z_{3}+z_{4})+2Nk_{\text{Bloch}}(z_{\text{u}}-z_{\text{d}})$ (35)
+$\Delta\phi_{\text{l}}=2k_{\text{Bragg}}(z_{1}-z_{2}-z_{3}+z_{4})+2Nk_{\text{Bloch}}(z_{\rm u}-z_{\rm d})$ (35)
 Let us notice the following relations between particular heights:
-$z_{2}-z_{1}=v_{1}T-\frac{1}{2}gT^{2},\qquad z_{4}-z_{3}=(v_{3}-\Delta v)T-\frac{1}{2}gT^{2},\qquad z_{\text{u}}-z_{\text{d}}=\Delta z,$ (36)
+$z_{2}-z_{1}=v_{1}T-\frac{1}{2}gT^{2},\qquad z_{4}-z_{3}=(v_{3}-\Delta v)T-\frac{1}{2}gT^{2},\qquad z_{\rm u}-z_{\rm d}=\Delta z,$ (36)
 where $v_{1}$ and $v_{3}$ are given by (28) and (30). Hence, we can write the total laser phase difference as
-$\Delta\phi_{\text{l}}=\frac{m}{\hbar}\left[-g\Delta z(T+2T^{\prime}+\tau)+g\Delta zN\tau_{\text{B}}\right].$ (37)
+$\Delta\phi_{\text{l}}=\frac{m}{\hbar}\left[-g\Delta z(T+2T^{\prime}+\tau)+g\Delta zN\tau_{\rm B}\right].$ (37)
 The total phase difference $\Delta\phi$ between the upper and the lower (ground state) trajectories is calculated by summing up the total propagation phase difference and the total laser phase difference.
@@
 $\Delta\phi=\Delta\phi_{\text{p}}+\Delta\phi_{\text{l}}=-\frac{mg\Delta z}{\hbar}(T+2T^{\prime}+\tau).$ (38)
-Now, let us calculate the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state on the lower and upper trajectory, respectively.
+Now, let us calculate the additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired by the excited state on the lower and upper trajectory, respectively.
 We need to calculate the additional laser phase coming from the clock pulses and the correction to the propagation phase due to the mass-energy equivalence.
@@
 Note that this phase is independent of the position $z$ of the interaction (thus, it is the same for the lower and upper trajectory).
-Since the excited state trajectories absorb the photons at $t_{\text{i}}$ and emit at $t_{\text{f}}$, they acquire the laser phase $\omega(t_{\text{f}}-t_{\text{i}})=\omega T_{\text{B}}$.
+Since the excited state trajectories absorb the photons at $t_{\text{i}}$ and emit at $t_{\text{f}}$, they acquire the laser phase $\omega(t_{\text{f}}-t_{\text{i}})=\omega T_{\rm B}$.
 To calculate the propagation phase correction, we employ the perturbative method developed in [62] and integrate the mass-energy correction to the free-fall Lagrangian
@@
 over the ground state trajectory on the corresponding height.
-For the free-fall trajectory starting at $t_{\text{A}}$ and ending at $t_{\text{B}}$ the correction $\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})$ to the propagation phase reads (compare with (20))
-$\delta_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=-\frac{\delta m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right].$ (40)
-For a full single oscillation in the optical lattice, the propagation phase correction $\delta_{\text{p,1}}$ is given by
-$\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (41)
-where $z_{\text{B}}$ and $v_{\text{B}}$ are the atom’s position and velocity at the interaction point, and $\tau_{\text{B}}$ is the period of the oscillation.
-We assume that the total number $N$ of Bloch oscillations is large, and neglect the correction coming from the fact that the total time $T_{\text{B}}$ of Bloch oscillations is not necessarily equal to the integer multiple of Bloch periods (instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$; however, since $\tau\ll N\tau_{\text{B}}$, we will assume $\tau\approx 0$).
+For the free-fall trajectory starting at $t_{\text{a}}$ and ending at $t_{\rm B}$ the correction $\delta_{\text{p}}(t_{\text{a}}\to t_{\rm B})$ to the propagation phase reads (compare with (20))
+$\delta_{\text{p}}(t_{\text{a}}\to t_{\rm B})=-\frac{\delta m}{\hbar}\left[\left(c^{2}+gz_{\text{a}}-\frac{1}{2}v_{\text{a}}^{2}\right)(t_{\rm B}-t_{\text{a}})+v_{\text{a}}g(t_{\rm B}-t_{\text{a}})^{2}-\frac{1}{3}g^{2}(t_{\rm B}-t_{\text{a}})^{3}\right].$ (40)
+For a full single oscillation in the optical lattice starting at height $z_{\rm B}$ with velocity $v_{\rm B}$, the propagation phase correction $\delta_{\text{p,1}}$ is given by
+$\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left[-c^{2}-g\left(z_{\rm B}+\frac{v_{\rm B}^{2}}{3}\right)+\frac{v_{\rm B}^{2}}{6}\right]\tau_{\rm B}\equiv-\omega_{0}\tau_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right],$ (41)
+where $\tau_{\rm B}$ is the oscillation period, we have used the fact that $\delta m=\hbar\omega_{0}/c^{2}$, and introduced $\langle z\rangle$ and $\langle v^{2}\rangle$ for the mean height and velocity squared.
+We assume that the total number $N$ of Bloch oscillations is large, and neglect the correction coming from the fact that the total time $T_{\rm B}$ of Bloch oscillations is not necessarily equal to the integer multiple of Bloch periods (instead, $T_{\rm B}=N\tau_{\rm B}+\tau$ with $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$; however, since $\tau\ll N\tau_{\rm B}$, we will assume $\tau\approx 0$).
 Then, the total propagation phase correction is given by
-$\delta_{\text{p,tot}}=N\delta_{\text{p,1}}=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)N\tau_{\text{B}}\approx\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)T_{\text{B}}.$ (42)
-Without loss of generality, we can assume that the interaction height corresponding to the lower trajectory $z_{\text{B}}^{\text{(d)}}=0$, while the interaction height of the upper one equals $z_{\text{B}}^{\text{(u)}}=\Delta z$.
-We introduce the mean velocity (within a single oscillation) $\langle v^{2}\rangle\equiv v_{\text{B}}^{2}/3$, and note that $\delta m=\omega_{0}/c^{2}$.
+$\delta_{\text{p,tot}}\approx N\delta_{\text{p,1}}=-\omega_{0}N\tau_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right]\approx-\omega_{0}T_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right].$ (42)
+Without loss of generality, we can assume that the mean height corresponding to the lower trajectory $\langle z\rangle_{\rm d}=0$, while this of the upper one equals $\langle z\rangle_{\rm u}=\Delta z$.
 Combining the propagation phase with the laser phase corrections, we get
-$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}}.\end{split}$ (43)
+$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1+\frac{g\Delta z}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B}.\end{split}$ (43)
 Appendix D Non-perturbative calculations
@@
 The trajectory followed by both trajectories has been described above.
-We analyze only the time between the two first interactions on the ground state trajectory (at some arbitrary height $z_{\text{B}}$) and show that the phase difference between the ground and the excited state trajectories agrees with the perturbative calculations.
+We analyze only the time between the two first interactions on the ground state trajectory (at some arbitrary height $z_{\rm B}$) and show that the phase difference between the ground and the excited state trajectories agrees with the perturbative calculations.
 It is straightforward that it will agree with the perturbative calculations also at later times.
@@
 (this is just Eq.
-(20) with $t_{\text{A}}\to 0$ and $t_{\text{B}}\to t$) where $m^{\text{(g)}}=m$ and $m^{\text{(e)}}=m+\delta m$ are the masses of the atom in the ground and excited state, respectively, $v_{0}$ is the initial velocity, and $z_{0}$ is the initial height.
-For the ground state, we replace $v_{0}$ by $v_{\text{B}}$, while for the excited state by $v_{\text{B}}-\delta v_{\text{B}}$.
-The initial height for both states is the same and equal to $z_{\text{B}}$.
+(20) with $t_{\text{A}}\to 0$ and $t_{\rm B}\to t$) where $m^{\text{(g)}}=m$ and $m^{\text{(e)}}=m+\delta m$ are the masses of the atom in the ground and excited state, respectively, $v_{0}$ is the initial velocity, and $z_{0}$ is the initial height.
+For the ground state, we replace $v_{0}$ by $v_{\rm B}$, while for the excited state by $v_{\rm B}-\delta v_{\rm B}$.
+The initial height for both states is the same and equal to $z_{\rm B}$.
 The difference in propagation phases at time $t$ is then given by
-$\delta_{\text{p}}(t)=\phi_{\text{p}}^{\text{(e)}}(t)-\phi_{\text{p}}^{\text{(g)}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right]-\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t$ (46)
+$\delta_{\text{p}}(t)=\phi_{\text{p}}^{\text{(e)}}(t)-\phi_{\text{p}}^{\text{(g)}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)t-v_{\rm B}gt^{2}+\frac{1}{3}g^{2}t^{3}\right]-\frac{m}{\hbar}(v_{\rm B}-gt)\delta v_{\rm B}t$ (46)
 (recall that we neglect the terms higher-order than $1/c^{2}$).
-Let us notice that, since $\delta v_{\text{B}}t$ is just the separation $\Delta z(t)$ between the trajectories, the second term in the above formula is equal in value but with the opposite sign to the separation phase
-$\delta_{\text{s}}(t)=\frac{\bar{p}(t)\delta v_{\text{B}}t}{\hbar}=\frac{m}{\hbar}(v_{\text{B}}-gt)\delta v_{\text{B}}t,$ (47)
-where $\bar{p}(t)=m(v_{\text{B}}-gt)+\mathcal{O}(1/c^{2})$. Therefore, the total phase difference for $t<\tau_{\text{B}}-\delta v_{\text{B}}/g$ is given by
-$\delta_{\text{p}}(t)+\delta_{\text{s}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)t-v_{\text{B}}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (48)
+Let us notice that, since $\delta v_{\rm B}t$ is just the separation $\Delta z(t)$ between the trajectories, the second term in the above formula is equal in value but with the opposite sign to the separation phase
+$\delta_{\text{s}}(t)=\frac{\bar{p}(t)\delta v_{\rm B}t}{\hbar}=\frac{m}{\hbar}(v_{\rm B}-gt)\delta v_{\rm B}t,$ (47)
+where $\bar{p}(t)=m(v_{\rm B}-gt)+\mathcal{O}(1/c^{2})$. Therefore, the total phase difference for $t<\tau_{\rm B}-\delta v_{\rm B}/g$ is given by
+$\delta_{\text{p}}(t)+\delta_{\text{s}}(t)=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)t-v_{\rm B}gt^{2}+\frac{1}{3}g^{2}t^{3}\right],$ (48)
 in agreement with the perturbative results (indeed, this is just the propagation phase correction calculated in the perturbative approach).
-At the time $\tau_{\text{B}}-\delta v_{\text{B}}/g$ the interaction of the atom on the excited trajectory with the laser occurs, and it acquires a laser phase
-$\phi_{\text{l}}^{\text{(e)}}=\frac{m}{\hbar}2v_{\text{B}}(z_{\text{B}}-\delta z_{\text{B}})=\phi_{\text{l}}^{\text{(g)}}-\frac{m}{\hbar}\delta v_{\text{B}}\tau_{\text{B}}+\mathcal{O}(\delta m^{2}),$ (49)
-where $\phi_{\text{l}}^{\text{(g)}}$ is the laser phase acquired by the ground state trajectory at time $\tau_{\text{B}}$.
+At the time $\tau_{\rm B}-\delta v_{\rm B}/g$ the interaction of the atom on the excited trajectory with the laser occurs, and it acquires a laser phase
+$\phi_{\text{l}}^{\text{(e)}}=\frac{m}{\hbar}2v_{\rm B}(z_{\rm B}-\delta z_{\rm B})=\phi_{\text{l}}^{\text{(g)}}-\frac{m}{\hbar}\delta v_{\rm B}\tau_{\rm B}+\mathcal{O}(\delta m^{2}),$ (49)
+where $\phi_{\text{l}}^{\text{(g)}}$ is the laser phase acquired by the ground state trajectory at time $\tau_{\rm B}$.
 Let us notice that right before the second interaction of the excited state the separation phase is equal to
-$\delta_{\text{s}}(\tau_{\text{B}}-\Delta v/g-\varepsilon)=-\frac{m}{\hbar}v_{\text{B}}\delta v_{\text{B}}t,$ (50)
+$\delta_{\text{s}}(\tau_{\rm B}-\Delta v/g-\varepsilon)=-\frac{m}{\hbar}v_{\rm B}\delta v_{\rm B}t,$ (50)
 while right after the interaction it is $\sim\mathcal{O}(1/c^{4})$, because the mean momentum is then $\sim\mathcal{O}(1/c^{2})$.
-This situation (mean momentum $\sim\mathcal{O}(1/c^{2})$ and thus negligible value of separation phase) persists until the second interaction on the ground state trajectory, at which time both trajectories are at the same height, but the excited one has the velocity $v_{\text{B}}-2\delta v_{\text{B}}$.
-At this point, we can repeat the whole analysis changing everywhere $\delta v_{\text{B}}\to 2\delta v_{\text{B}}$.
+This situation (mean momentum $\sim\mathcal{O}(1/c^{2})$ and thus negligible value of the separation phase) persists until the second interaction on the ground state trajectory, at which time both trajectories are at the same height, but the excited one has the velocity $v_{\rm B}-2\delta v_{\rm B}$.
+At this point, we can repeat the entire analysis changing everywhere $\delta v_{\rm B}\to 2\delta v_{\rm B}$.
 Since the propagation phase difference acquired between the second interaction on the excited and on the ground trajectories is also $\sim\mathcal{O}(\delta m^{2})$, the total phase difference after the first oscillation is equal to
-$\delta=\delta_{\text{p}}+\delta_{\text{s}}+\delta_{\text{l}}=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\text{B}}^{2}-gz_{\text{B}}\right)\tau_{\text{B}}-v_{\text{B}}g\tau_{\text{B}}^{2}+\frac{1}{3}g^{2}\tau_{\text{B}}^{3}\right]=\frac{\delta m}{\hbar}\left(-c^{2}-\frac{v_{\text{B}}^{2}}{6}-gz_{\text{B}}\right)\tau_{\text{B}},$ (51)
+$\delta=\delta_{\text{p}}+\delta_{\text{s}}+\delta_{\text{l}}=\frac{\delta m}{\hbar}\left[\left(-c^{2}+\frac{1}{2}v_{\rm B}^{2}-gz_{\rm B}\right)\tau_{\rm B}-v_{\rm B}g\tau_{\rm B}^{2}+\frac{1}{3}g^{2}\tau_{\rm B}^{3}\right]=\frac{\delta m}{\hbar}\left[-c^{2}-g\left(z_{\rm B}+\frac{v_{\rm B}^{2}}{3}\right)+\frac{v_{\rm B}^{2}}{6}\right]\tau_{\rm B},$ (51)
 which is equal to the propagation phase correction calculated for a single oscillation in the perturbative approach (compare with Eq. (41)).
+Appendix E Bloch oscillations — the usual description
+In our analysis, we have employed a simplified model of Bloch oscillations based on [61, 19], where the oscillating atom follows a bouncing trajectory in a gravitational field (see Fig.
+2).
+However, this is not the most common description of Bloch oscillations — usually one assumes that the dynamics of the atom is set by a periodic potential (optical lattice) and a linear correction (gravitational field), which leads to the oscillatory motion [68, 69].
+The full Hamiltonian of the system involving the leading relativistic corrections reads
+[…]
+where $V_{0}$ is the depth of the optical lattice potential, $k_{\text{Bloch}}$ is the wave vector of the lattice photons, $|\ket{e}\bra{e}$ is a projector on the excited state of the atom, and
+[…]
+Writing Eq.
+(52) we have assumed that the lattice operates at the “magic” wavelength [58, 72], which means that the harmonic potential experienced by the ground and excited state atoms is the same.
+Let us notice that the frequency of the light as perceived by the atom depends on its position and velocity (due to the relativistic Doppler shifts).
+Therefore, in principle, it can only be at the “magic” wavelength for an atom at some specific height and velocity.
+The atom at some other height, or moving with a different velocity, will perceive the light as shifted from this wavelength, leading to a dependence of the optical lattice potential on the center-of-mass degrees of freedom.
+Note that this is exactly analogous to the situation in optical lattice clocks, where the atoms are not perfectly static and are placed at different heights.
+However, it was shown in [58] that around the “magic” wavelength the potential changes very slowly with the frequency of the lattice laser, and the corrections due to Doppler shifts can be made negligible.
+Consequently, we shall omit them from further consideration.
+[…]
+$\delta_{\rm d}^{(0)}=\delta_{\rm u}^{(0)}=(\omega-\omega_{0})T_{\rm B}.$ (54)
+The first-order (in $1/c^{2}$) correction to the phases results from the correction to the eigenenergies and is given by
+$\delta_{i}^{(1)}=-\frac{1}{\hbar}\left(\left\langle\hat{H}^{(1)}\right\rangle_{e,i}-\left\langle\hat{H}^{(1)}\right\rangle_{g,i}\right)T_{\rm B},$ (55)
+where the expectation values are evaluated on unperturbed states. In the ground state, the expectation value of $\hat{H}^{(1)}$ is zero irrespectively of the path, and in the excited state we have
+$\left\langle\hat{H}^{(1)}\right\rangle_{e,i}=\frac{\hbar\omega_{0}}{c^{2}}\left(g\langle\hat{z}\rangle_{i}-\frac{\langle\hat{p}^{2}\rangle_{i}}{2m^{2}}\right).$ (56)
+This leads to the total phase shift of the form:
+$\delta_{i}=\delta_{i}^{(0)}+\delta_{i}^{(1)}=\left[\omega-\omega_{0}\left(1+\frac{g\langle\hat{z}\rangle_{i}}{c^{2}}-\frac{\langle\hat{p}^{2}\rangle_{i}}{2m^{2}c^{2}}\right)\right]T_{\rm B},$ (57)
+which, upon identifications $\langle\hat{z}\rangle_{i}$ and $\langle\hat{p}^{2}\rangle_{i}=m^{2}\langle v^{2}\rangle_{i}$, takes the same form as derived in Appendix C.
+Note the different meaning of the brackets in (57) and those present e.g.
+in (43) — in the first case they indicate averaging over the quantum state, while in the latter they mean taking the average over the classical trajectory.
+The same results can be derived in the Lagrangian formalism using the method developed in [62].
+We assume that the atom follows a superposition of classical trajectories set by the (ground-state) Lagrangian
+$L_{g}=-mc^{2}+\frac{1}{2}mv^{2}-mgz-V_{0}\cos(2k_{\text{Bloch}}z),$ (58)
+corresponding to the Hamiltonian (52). Let us consider (following the reasoning from Appendix C) the dominant correction to the Lagrangian of the excited state atoms, i.e.,
+$\delta L=-\delta m\left(c^{2}+gz-v^{2}/2\right).$ (59)
+To calculate the phases $\delta_{\rm d}$ and $\delta_{\rm u}$ ($\Delta\phi$ does not depend on the model of Bloch oscillations that we use) we integrate the Lagrangian correction $\delta L$ over the trajectory $z(t)$ followed by the atom (assuming that the trajectory is the same for both internal states).
+This gives (recall that $\delta m=\hbar\omega_{0}/c^{2}$)
+$\frac{1}{\hbar}\int_{0}^{T_{\rm B}}\mathrm{d}t\delta L=-\omega_{0}T_{\rm B}\left[1+\frac{g}{c^{2}}\left(\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tz(t)\right)-\frac{1}{2c^{2}}\left(\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tv^{2}(t)\right)\right]\equiv-\omega_{0}T_{\rm B}\left[1+\frac{g\langle z\rangle}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right],$ (60)
+where we have introduced
+$\langle z\rangle\equiv\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tz(t),\qquad\langle v^{2}\rangle\equiv\frac{1}{T_{\rm B}}\int_{0}^{T_{\rm B}}\mathrm{d}tv^{2}(t).$ (61)
+Assuming that the mean height of the lower trajectory is set to zero and that of the upper one to $\Delta z$, that the mean squared velocity is the same on both heights and adding a laser phase $\omega T_{\rm B}$ coming from the interaction with the clock pulse, we get
+$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1+\frac{g\Delta z}{c^{2}}-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\end{split}$ (62)
+which is the same as derived in Appendix C.
+It is worth noting that the method developed in [62] was derived for quadratic Lagrangians, while Lagrangian (58) is nonquadratic.
+However, the compatibility with the result obtained in the Hamiltonian formalism suggests that the method from [62] is still applicable in our case.